### v1

# GranQ: Granular Zero-Shot Quantization with Unified Layer-Channel Awareness

Abstract
Zero-shot quantization (ZSQ) enables neural network compression without training data, which is crucial in restricted data access environments. However, existing ZSQ methods suffer from significant activation loss in low-bit environments owing to their coarse-grained scaling strategy. To address this issue, we propose GranQ, a novel ZSQ approach that leverages layer-channel awareness to minimize the quantization error. Unlike conventional layer- or channel-wise quantization, GranQ dynamically adjusts quantization granularity by considering both layer- and channel-level activation distributions. This enables fine-grained quantization while minimizing activation distortion. Additionally, we introduce vectorized activation quantization, which enables efficient parallel computation and reduces computational overhead while preserving accuracy. GranQ achieves superior performance compared with those of state-of-the-art ZSQ methods that employ quantization-aware training. With these findings, we anticipate that GranQ will inspire novel research directions beyond conventional ZSQ approaches focused on data generation and model training. The official code will be released upon completion of the review.
1 Introduction
Figure 1: Comparison of quantization strategies. Existing methods apply either per-tensor scaling with layer-wise quantization or per-channel scaling. GranQ combines per-channel scaling with per-layer iteration for more granular quantization.
Neural network compression techniques have been extensively studied to make large-scale deep learning (DL) models practically deployable. In particular, reducing model size while minimizing performance degradation is crucial for utilizing DL on edge devices (e.g., mobile phones, embedded systems, and drones). Major approaches to model compression [7, 12] include quantization [24, 30, 17], pruning [28, 16, 33, 6], knowledge distillation [21, 20, 40], and neural architecture search [45, 13]. Among these, quantization has emerged as the most actively studied technique. It serves as an effective compression method by reducing unnecessary representational ranges in the model while maintaining its performance. However, a major limitation of quantization is the need for fine-tuning or calibration to maintain the performance of the full-precision (FP) model [17]. To address this limitation, zero-shot quantization (ZSQ) (i.e., data-free quantization) has been introduced [34, 41], which allows compressed models to maintain their performance without requiring access to the original training data. Since the introduction of ZeroQ [4], studies on ZSQ have advanced in two main directions. The first direction focuses on data generation, where synthetic data are created from the FP model. The second direction focuses on effectively applying the activation distributions of the synthetic data to the quantized (Q) model. This second direction is further divided into post-training quantization (PTQ), which calibrates the activation distributions, and quantization-aware training (QAT), which fine-tunes the Q model directly. We categorize existing ZSQ methods based on these directions, as summarized in Table 1.
Data Generation
(PTQ, QAT)
Calibration
Fine-tuning
ZeroQ (CVPR 20) [4]
GDFQ (ECCV 20) [41]
DSG (CVPR 21) [43]
MixMix (CVPR 21) [31]
Qimera (NeurIPS 21) [9]
IntraQ (CVPR 22) [44]
GENIE (CVPR 23) [25]
AdaDFQ (CVPR 23) [36]
Casual-DFQ (ICCV 23) [39]
TexQ (NeurIPS 23) [5]
RIS (AAAI 24) [1]
GenQ (ECCV 24) [32]
SQuant (ICLR 22) [18]
UDFC (ICCV 23) [2]
AIT (CVPR 22) [10]
PLF (CVPR 24) [14]
SynQ (ICLR 25) [26]
GranQ (Ours)
Table 1: Categorizing zero-shot quantization algorithms. PTQ denotes post-training quantization, and QAT denotes quantization-aware training. Data generation methods typically support both PTQ and QAT, whereas violet-marked methods utilize fine-tuning for data generation and support only QAT.
First, data generation studies focus on generating high-quality data to effectively train the Q model. A common approach uses batch normalization statistics from the FP model [41]. Recent studies focused on generating data adaptable to various bit-widths [36] and using diverse generative models such as stable diffusion guided by text prompts [32]. Meanwhile, calibration (PTQ) studies aim to minimize the quantization error by calibrating the Q model without additional training [18, 2]. Finally, fine-tuning (QAT) studies focus on transferring key information from the FP model to the Q model through knowledge distillation [10, 14, 23, 26].
However, despite extensive studies, severe performance degradation in low-bit quantization remains unresolved. To address this issue, we performed an in-depth analysis of the ZSQ process, focusing on why low-bitwidth settings still suffer from performance loss even after QAT fine-tuning. Our findings reveal that quantization errors mainly stem from the loss of activation values instead of data quality or training methods. Notably, we discovered that layer-wise (per-tensor) quantization fails to accurately preserve activations in ZSQ, leading to coarse and inaccurate representations.
Based on this analysis, we introduce GranQ, a novel ZSQ approach. GranQ independently handles large activations in specific layers and channels by narrowing the range between the maximum and minimum values within each channel (Figure 1). This dynamic adjustment minimizes activation loss and preserves the original activation values by reducing quantization errors, as shown in Figure 2. The proposed method effectively handles activation loss in low-bit quantization and achieves state-of-the-art (SOTA) performance in QAT settings on the CIFAR and ImageNet datasets. Furthermore, we implement vectorized operations for granular quantization with minimal computational overhead, enabling efficient quantization on GPUs. Our contributions can be summarized as follows:
  • We perform a comprehensive analysis of quantization errors in low-bit ZSQ settings. Our findings reveal that conventional activation quantization methods (e.g., layer- and channel-wise) lead to significant activation loss. This issue stems from coarse-grained activation quantization, which fails to preserve fine-grained details.
  • We propose GranQ, a novel method that supports granular quantization and maintains computational efficiency. We introduce a fine-grained quantization mechanism, which considers every channel in each layer, to reduce activation loss. Additionally, we mitigate computational overhead using vectorized activation quantization. To the best of our knowledge, our approach is the first to address the ZSQ problem.
  • We achieve SOTA performance over existing ZSQ methods through extensive evaluation. Specifically, on the CIFAR-100 dataset, in the 3-bit quantization setting, GranQ achieves an accuracy of 62.73%, improving by 5.45% over the latest method on ResNet-20. Furthermore, on the CIFAR-10 dataset in the 5-bit quantization setting, GranQ achieves an accuracy of 94.06%, slightly exceeding the FP model performance by 0.17%.
(a) Layer Quantization
(b) GranQ (Ours)
Figure 2: Comparison between (a) layer-wise quantization and (b) the proposed GranQ method on the CIFAR-10 dataset from AdaDFQ[36]. We visualize the activation of the first layer in ResNet-20 for the first image of the first batch. $w$ and $a$ denote weight and activation, respectively. Each sub-figure presents the activation graph for 32-bit FP (left) and 3-bit quantization (right). Although layer-wise quantization significantly distorts the original activation after quantization, GranQ effectively preserves it with minimal distortion.
2 Related Work
2.1 Quantization
Quantization can reduce the representational range of pre-trained deep neural networks (DNNs) and minimize memory usage. Quantization methods are categorized into PTQ and QAT based on the timing of the range reduction.
PTQ quantizes fully trained DNNs without any further training [17, 3], offering low computational cost and simple implementation. However, because PTQ directly applies quantization to the FP model, it struggles to properly adjust the range during quantization. To overcome this, calibration is used to select optimal quantization ranges by determining scaling factors. In contrast, QAT fine-tunes the Q model with quantization operations integrated into the training process, enabling direct optimization of activations [17]. In particular, it employs the straight-through estimator [42] to estimate gradients.
However, both PTQ and QAT have the limitation of requiring access to data. PTQ needs data for calibration, whereas QAT requires it for fine-tuning. In practice, obtaining training data is often challenging. To address this problem, ZSQ has been proposed, which performs quantization without using original data [41].
2.2 Zero-shot Quantization
As summarized in Table 1, ZSQ was initially introduced by ZeroQ [4], leading to the development of various ZSQ algorithms. These studies have explored diverse approaches to improve quantization performance under data-free settings. Recent studies have focused on generating high-quality synthetic data [25, 32] and developing more effective methods to utilize them [10, 23, 26]. However, these approaches are still limited in preserving activations in low-bit settings. Therefore, a new quantization method is required to effectively minimize activation loss. In this study, we propose a novel approach to address this challenge.
3 Preliminaries and Problem Definition
In this section, we provide a new definition of the existing ZSQ problem and introduce the preliminaries. The detailed mathematical notations are listed in Table LABEL:tab6 in the supplementary material.
3.1 Activation Quantization
Activation quantization reduces the precision of intermediate activation values in neural networks by converting them into low-bitwidth representations [8]. This technique is often used alongside weight quantization to further compress the model. Among various activation quantization methods, linear quantization is the most widely used. It uses a normalized scaling factor to adjust the activation values, and subsequently converts them into integers within a fixed bit range [24, 15]. The overall process typically involves three stages: forward quantization, scaling parameter computation, and dequantization [17].
3.1.1 Forward Quantization
The forward quantization process involves mapping continuous floating-point values to an integer space. This process utilizes the quantization operator $Q$, and the equation is defined as follows:
$x_{q}=\mathcal{Q}(x,s,z)=\left\lfloor\frac{x}{s}+z\right\rceil$ (1)
[…]
3.1.2 Scaling Parameter Computation
In activation quantization, setting the scaling factor $s$ is crucial for mapping various activation distributions to a fixed integer range. In this study, we employ the asymmetric quantization method, which is the most commonly used for activation quantization [29]. We calculate the scaling factor $s$ and the zero-point $z$ as follows:
$s=\frac{x_{\max}-x_{\min}}{2^{b}-1},\quad z=\left\lfloor\frac{-x_{\min}}{s}\right\rceil$ (2)
Here, $x_{\max}$, $x_{\min}$ and $b$ represent the maximum and minimum values of the activation, and the number of quantization bits, respectively. The zero-point $z$ represents the offset used to adjust activation values. $x_{\max}$ and $x_{\min}$ determine the range of the activation distribution, which directly affects the computation of scaling factor $s$. Specifically, when $x_{\max}-x_{\min}$ is large, the scaling factor $s$ increases. This leads to a larger quantization step size, making it harder to preserve fine-grained activations. Conversely, when $x_{\max}-x_{\min}$ is small, the scaling factor $s$ also decreases. This allows for finer value distinctions and enhances the quality of representation.
3.1.3 Dequantization
The dequantization process converts quantized integer values back into their real-valued representation, as shown in the following equation:
$x_{\text{deq}}=s(x_{q}-z)$ (3)
In this process, the zero-point $z$ is subtracted to reverse the forward quantization, and the result is multiplied by the scaling factor $s$ to obtain the final value $x_{deq}$. The dequantization process restores quantized activation values to their original real-valued range. It is essential during model inference for converting final predictions into real values. This process is essential for converting the final prediction into real values during model inference.
Figure 3: Visualization of activation magnitudes for the first image in the first batch using ResNet-20 on CIFAR-10. The left column illustrates activations from the FP model, whereas the right column presents activations after layer-wise quantization. Significant distortions are observed in certain layers, particularly in the final layer.
3.2 Problem Definition
Existing activation quantization methods are designed based on layer-wise quantization to minimize computational cost. However, fixed scaling factors struggle to handle varying activations. This issue becomes more severe under low-bitwidth settings in ZSQ environments. The following outlines the primary challenges faced in ZSQ under low-bitwidth settings.
(Problem 1) Coarse-grained activation representation
In conventional activation quantization approaches, the scaling factor $s$ is fixed per layer or channel, mapping all activations to the same range. However, in low-bitwidth settings, this approach struggles to capture fine-grained activation variations from generated data.
(Problem 2) Severe quantization error in certain layers
During quantization, certain layers induce severe quantization errors, leading to significant activation distortions. These distortions substantially differ from the original distribution and can directly impact the performance of the model.
4 Observation and Methodology
We observe the causes of activation distortion during quantization and analyze how to mitigate them. Based on this, we introduce the proposed method known as GranQ.
4.1 Observation
ResNet-20 Layer-Q Channel-Q GranQ
$Avg.\frac{X\cdot Q}{\|X\|_{2}\|Q\|_{2}}(\uparrow)$
$Avg.\frac{\|X-Q\|_{2}}{\|X\|_{2}}(\downarrow)$
0.5111
0.3129
0.5602
0.2528
0.6835
0.1063
Table 2: Comparative analysis of averaged cosine similarity and relative error over all layers and batches. Higher cosine similarity ($\uparrow$) indicates better retention of activation information, whereas lower relative error ($\downarrow$) signifies reduced quantization distortion. Layer-Q and Channel-Q denote layer- and channel-wise quantization, respectively, whereas GranQ integrates both for improved quantization.
Layer-wise quantization effectively captures the activation distribution of each layer. However, it becomes inefficient when activation distributions within a layer exhibit significant variation along the channel axis. Particularly in ZSQ, as shown in the FP graph in Figure 2(a), the activation distributions can differ substantially across channels even within the same layer. This implies that limited representation bits are forced to capture a wide range of activation values, making fine-grained quantization difficult. It consistently appears across all layers. We analyzed this phenomenon occurring across all layers and provided a summary in Table 2. In Table 2, we present the cosine similarity $\frac{X\cdot Q}{|X|_{2}|Q|_{2}}$, along with the relative error $\frac{|X-Q|_{2}}{|X|_{2}}$. From Table 2, we identify that both layer- and channel-wise quantization fail to preserve the original activation effectively. This indicates that layer- and channel-wise quantization, which independently process activations at the layer or channel level, are not suitable for ZSQ. In contrast, GranQ, which considers both layers and channels simultaneously, achieves 1.34$\times$ higher similarity and 2.94$\times$ lower relative error compared with those of layer-wise quantization, demonstrating its effectiveness in preserving activations.
Furthermore, Figure 3 provides a visual representation of the effect of quantization on activation values across layers. In layer-quantizaiton, although the overall activations generally reflect the FP activations to some extent, certain layers exhibit fine activation values being mapped to discrete values, leading to coarse representations. These coarse representations have been identified as a major factor leading to a decline in the overall expressiveness and inference performance of the model. Further analysis of Figure 3 is provided in the supplementary material, in Figures LABEL:sup_fig3 and LABEL:sup_fig1.
4.2 Granular Quantization
Based on our analysis, we propose GranQ (granular layer-channel quantization), a fine-grained ZSQ approach designed to minimize computational overhead, as shown in Figure 4. GranQ combines layer and channel information using layer-channel scaling and improves efficiency through vectorized quantization.
4.2.1 Layer-Channel Scaling
We propose layer-channel scaling, which enables GranQ to assign individual scaling factors to each channel, enhancing quantization precision. This approach reduces the impact of outlier activations in specific channels on the quantization process. The layer-channel scaling is defined as follows:
$\vec{x}_{\min}$ $=\min_{h,w}(x_{l}),\quad\vec{x}_{\max}=\max_{h,w}(x_{l})$ (4)
where $\text{where }x_{l}\in\mathbb{R}^{C\times H\times W}$
[…]
Dataset
Model
(FP 32)
Bits
GDFQ
(ECCV 20)
ARC+AIT
(CVPR 22)
AdaDFQ
(CVPR 23)
TexQ
(NeurIPS 24)
AIT+RIS
(AAAI 24)
PLF
(CVPR 24)
GenQ
(ECCV 24)
AKT
(SAC 25)
SynQ
(ICLR 25)
GranQ
[…]
Table 3: Accuracy Evaluation of QAT Methods for ZSQ. $w$ and $a$ represent weight and activation, respectively. Bold values indicate the best accuracy, and underlined values denote the second-best accuracy. $\dagger$ indicates our re-implementation with the official code.
[…]
4.2.2 Vectorized Quantization
In the layer-channel scaling stage, the activation tensor of each layer is vectorized along the channel dimension. In the subsequent vectorized quantization stage, both the scaling factor $\vec{s}$ and zero-point $\vec{z}$ are also vectorized and computed as shown in Equation 5. The activation values go through quantization and dequantization (Equation 6), with all operations running in parallel across channels. During computation, scaling factor $s$ and zero-point $z$ of each channel are broadcasted to the activation values and applied element-wise.
$\vec{s}=\frac{\vec{x}_{\max}-\vec{x}_{\min}}{2^{b}-1},\quad\vec{z}=\left\lfloor-\frac{\vec{x}_{\min}}{\vec{s}}\right\rceil$ (5)
$x_{q,l}=\left\lfloor\frac{x}{\vec{s}}+\vec{z}\right\rceil,x_{\text{deq},l}=\vec{s}(x_{q,l}-\vec{z})$ (6)
This process can be applied to every layer in the network, enabling the vectorized quantization of all activation tensors. A detailed computational process is provided in Algorithm LABEL:alg1, available in the supplementary material.
5 Experiments
In this section, we thoroughly evaluate the effectiveness of GranQ. Experiments are conducted on diverse benchmark datasets, with the performance compared with those of existing ZSQ methods.
5.1 Experimental Setup and Details
The experiments were conducted using widely adopted ZSQ evaluation datasets, including CIFAR-10, CIFAR-100 [27], and ImageNet (ILSVRC 2012) [11] validation datasets. For the CIFAR datasets, ResNet-20 [19] was used as the quantization model, whereas ResNet-18 [19], ResNet-50 [19], and MobileNetV2 [38] were employed for ImageNet. All experiments were conducted using the SGD optimizer [37] with a momentum of 0.9 and weight decay of 1e-4. The CIFAR-10 and CIFAR-100 experiments were each conducted for 200 epochs, with batch sizes of 16 and 200, respectively. For ImageNet, we trained for 400 epochs with a batch size of 16. The initial learning rate was set to 1e-4 for CIFAR-10 and CIFAR-100, and 1e-5 for ImageNet, with multi-step learning rate decay applied. The decay steps were set to 100, 200, and 300 epochs for CIFAR, and at 350 and 400 epochs for ImageNet, with a decay rate of 0.1. We compared our method with existing ZSQ methods [41, 10, 36, 5, 1, 14, 32, 23, 26]. For data generation, we followed the AdaDFQ [36] approach based on ACGAN [35]. Further implementation details are provided in the supplementary material, Section LABEL:sec:Details_of_Experimental_Setup. In our ablation study, layer-wise quantization was applied to all layers containing activation functions, while channel-wise quantization was performed per channel at the batch level.
CIFAR-100 ResNet-20 (70.33%)
Method GDFQ Qimera+AIT AdaDFQ AdaDFQ+AKT
Baseline +GranQ Baseline +GranQ Baseline +GranQ Baseline +GranQ
3w3a 47.61 59.04+11.43 45.70 60.42+14.72 52.74 62.73+9.99 54.68 62.01+7.33
4w4a 63.39 66.97+3.58 65.80 68.08+2.28 66.81 68.79+1.98 66.94 68.77+1.83
5w5a 66.12 68.96+2.84 69.26 70.14+0.88 69.93 70.05+0.12 69.75 70.21+0.46
[…]
Layer-Q Channel-Q
Layer-Channel
Layer-Channel
(Vectorization, GranQ)
Accuracy (%)
49.98 (±$\pm 0.24$)
51.26 (±$\pm 0.28$)
62.68 (±$\pm 0.11$)
62.73 (±$\pm 0.17$)
Table 5: Comparative analysis of four quantization approaches on ResNet-20 accuracy. The experiments were conducted using the AdaDFQ method on CIFAR-100 with 3-bit quantization. Results are averaged over three repetitions with standard deviation.
5.2 Performance Evaluation
We evaluated the performance of GranQ against SOTA ZSQ methods, with results summarized in Table 3. All comparison experiments were conducted under 3, 4, and 5-bit quantization settings.
CIFAR-10/100. As indicated in Table 3, GranQ consistently achieved the highest accuracy across all bitwidths. For CIFAR-10, it attained accuracies of 94.06% (5-bit), 93.52% (4-bit), and 91.37% (3-bit). For CIFAR-100, the results were 70.05% (5-bit), 68.79% (4-bit), and 62.73% (3-bit). Notably, GranQ outperformed SynQ [26], the previous SOTA, by +5.45% in the CIFAR-100 3-bit setting. This result demonstrates its ability to effectively overcome the limitations of conventional low-bitwidth quantization techniques. Overall, GranQ consistently outperformed existing methods across all bitwidths in both CIFAR-10 and CIFAR-100, with particularly strong improvements in the 3-bit quantization setting. Remarkably, in the CIFAR-10 5-bit setting, GranQ even surpassed the performance of the FP model. These results suggest that we can effectively apply GranQ to small- and medium-scale datasets.
ImageNet. In the ImageNet experiments, GranQ achieved competitive performance across various bitwidths. Specifically, for ResNet-18, it attained top accuracies of 70.39% (4-bit) and 71.31% (5-bit). For ResNet-50, it achieved the highest accuracies of 76.63% (4-bit) and 71.31% (5-bit). Additionally, in the MobileNetV2 setting, GranQ achieved SOTA performance across all bitwidths (3, 4, and 5-bit). In the 3-bit setting, GranQ achieved the second-best accuracies, with 64.41% on ResNet-18 and 70.76% on ResNet-50. These results are -3.77% and -3.23% lower than those of GenQ, respectively. However, GenQ uses a pre-trained diffusion model [22] and prompts during data generation, leading to slower speeds compared with those of typical GAN-based methods [41]. In contrast, GranQ maintains the speed advantage of GAN-based generation while achieving high performance through fine-grained adjustments.
5.3 Ablation Study
5.3.1 Effectiveness Evaluation
GranQ consistently demonstrates performance improvements when applied to various ZSQ methods. As summarized in Table 4, GranQ achieves steady performance improvements when integrated with existing SOTA ZSQ techniques.
First, we analyzed the impact of GranQ on data synthesis-based quantization methods, specifically GDFQ [41] and AdaDFQ [36]. The integration of GranQ into GDFQ [41] led to a notable improvement, with accuracy rising from 47.61% to 59.04% (+11.43%) in the 3-bit setting and from 63.39% to 66.97% (+3.58%) in the 4-bit setting. Similarly, AdaDFQ [36] exhibited improvements of +9.99% in the 3-bit setting (from 52.74% to 62.73%) and +1.98% in the 4-bit setting (from 66.81% to 68.79%). These results suggest that GranQ effectively reduces quantization errors when combined with data synthesis-based quantization methods, leading to enhanced model performance.
Furthermore, GranQ also exhibits significant improvements in methods focused on training quantized models, such as Qimera+AIT [10] and AdaDFQ+AKT [23]. For Qimera+AIT [10], the 3-bit accuracy increased from 45.70% to 60.42% (+14.72%). Similarly, for AdaDFQ+AKT [23], the accuracy improved from 54.68% to 62.01% (+7.33%). These findings demonstrate that GranQ is not only effective in data synthesis-based methods but also enhances performance during the model training process.
Figure 5: Ablation study on the latency of the quantization process in ResNet-20 across different batch sizes. The experiments were conducted using the AdaDFQ method on CIFAR-100 with a 3-bit quantization setting.
5.3.2 Efficiency Evaluation
In general, performing precise quantization that considers both layers and channels leads to an exponential increase in computational cost. However, GranQ is designed to minimize this overhead while maintaining high accuracy by leveraging vectorized operations. To validate this, we measured the latency across various batch sizes (16, 32, 64, 128, and 200). Figure 5 compares four methods.
The first two methods are conventional quantization techniques: layer- and channel-wise. The third is scalar-based iterative, which considers all channels in each layer but performs multiple iterations. The fourth method is GranQ, which integrates both layer and channel information while employing vectorized quantization. Although scalar-based iterative quantization performs the same operations as GranQ, it relies on repeated scalar computations, as illustrated in Equations 1, 2, and 3. In contrast, GranQ enhances efficiency through parallelized computations (Algorithm LABEL:alg1 in supplementary material).
A key observation from Figure 5 is that GranQ induces only a slight increase in latency compared with layer- and channel-wise quantization. Moreover, as summarized in Table 5, GranQ achieves significantly higher accuracy than traditional quantization methods. Even when compared with the less efficient iteration-based layer-channel quantization, GranQ maintains nearly identical accuracy. Additionally, as summarized in Table 2, GranQ substantially reduces quantization error and effectively preserves activations. These findings confirm that GranQ achieves fine-grained quantization while benefiting from parallel computation.
6 Discussion
6.1 Why GranQ is Effective?
The most common activation quantization approaches are known for their low computational cost and fast processing speeds. However, they often fail to accurately capture activation distributions, resulting in substantial quantization errors (Figure 3). In layer-wise quantization, representing the entire activation range with a single scaling factor makes it difficult to reflect fine-grained distribution changes. Similarly, channel-wise quantization encounters challenges owing to significant range differences in activations across layers for the same channel, leading to increased information loss (Section LABEL:sec:extraanalysis in the supplementary material). These limitations become more pronounced in low-bit settings, where higher quantization errors result in performance degradation.
To overcome these issues, GranQ introduces the layer-channel scaling method, which combines layer- and channel-wise quantization. Unlike conventional methods, GranQ considers the activation distribution of each channel individually while also incorporating layer information, enabling more precise quantization. Furthermore, although conventional quantization relies on scalar operations with a single scaling factor, GranQ utilizes multiple scaling factors and performs vectorized computations. This approach leverages GPU resources efficiently, minimizing computational overhead while achieving effective quantization.
Experimental results demonstrate that GranQ outperforms traditional layer- and channel-wise quantization methods, achieving higher cosine similarity and lower relative error (Table 2). This demonstrates that GranQ accurately preserves the original activation distribution. Additionally, GranQ maintains computational efficiency in terms of latency and effectively balances the trade-off between computational cost and performance (Figure 5 and Table 5). Most notably, GranQ achieves SOTA performance compared with those of existing layer-wise quantization methods (Table 3), highlighting its efficiency and superiority over scalar iteration approaches.
6.2 Limitation and Future-work
As GranQ aims to precisely preserve activations during quantization, it partially depends on generated data quality. Consequently, studies on data generation remain highly important. The proposed method has the potential to achieve even better performance when applied to high-quality data generation strategies. In future work, we aim to improve the quantization performance by focusing on both the accurate generation of original activations and their effective preservation.
7 Conclusion
We present a comprehensive analysis of activation quantization errors in ZSQ and propose GranQ, a novel quantization method to effectively reduce these errors. Conventional layer- and channel-wise quantization methods rely on a single scaling factor, which often results in substantial activation loss. In contrast, GranQ employs a fine-grained quantization approach that simultaneously considers both layer and channel information, thereby reducing activation loss and enabling more precise quantization. Moreover, GranQ successfully maintains computational efficiency through vectorized quantization while performing detailed and fine-grained quantization.
Through extensive experiments, we demonstrate that GranQ achieves superior performance while preserving computational efficiency. Compared with various SOTA methods, GranQ demonstrates competitive performance and achieves notable improvements. Our study highlights the importance of activation preservation in ZSQ and proposes GranQ as an insightful direction to address this challenge.
Acknowledgement This research was supported by the National Research Foundation (NRF) funded by the Korean government (MSIT) (No. RS-2023-00229822).

### v3

@@
 3 Preliminaries and Problem Definition
-In this section, we provide a new definition of the existing ZSQ problem and introduce the preliminaries.
-The detailed mathematical notations are listed in Table LABEL:tab6 in the supplementary material.
+In this section, we provide a new definition of the existing ZSQ problem and introduce the preliminaries. The detailed mathematical notations are listed in Table 6 in the supplementary material.
 3.1 Activation Quantization
@@
 These coarse representations have been identified as a major factor leading to a decline in the overall expressiveness and inference performance of the model.
-Further analysis of Figure 3 is provided in the supplementary material, in Figures LABEL:sup_fig3 and LABEL:sup_fig1.
+Further analysis of Figure 3 is provided in the supplementary material, in Figures 6 and 7.
 4.2 Granular Quantization
@@
 This process can be applied to every layer in the network, enabling the vectorized quantization of all activation tensors.
-A detailed computational process is provided in Algorithm LABEL:alg1, available in the supplementary material.
+A detailed computational process is provided in Algorithm 1, available in the supplementary material.
 5 Experiments
@@
 For data generation, we followed the AdaDFQ [36] approach based on ACGAN [35].
-Further implementation details are provided in the supplementary material, Section LABEL:sec:Details_of_Experimental_Setup.
+Further implementation details are provided in the supplementary material, Section B.
 In our ablation study, layer-wise quantization was applied to all layers containing activation functions, while channel-wise quantization was performed per channel at the batch level.
@@
 Although scalar-based iterative quantization performs the same operations as GranQ, it relies on repeated scalar computations, as illustrated in Equations 1, 2, and 3.
-In contrast, GranQ enhances efficiency through parallelized computations (Algorithm LABEL:alg1 in supplementary material).
+In contrast, GranQ enhances efficiency through parallelized computations (Algorithm 1 in supplementary material).
 A key observation from Figure 5 is that GranQ induces only a slight increase in latency compared with layer- and channel-wise quantization.
@@
 In layer-wise quantization, representing the entire activation range with a single scaling factor makes it difficult to reflect fine-grained distribution changes.
-Similarly, channel-wise quantization encounters challenges owing to significant range differences in activations across layers for the same channel, leading to increased information loss (Section LABEL:sec:extraanalysis in the supplementary material).
+Similarly, channel-wise quantization encounters challenges owing to significant range differences in activations across layers for the same channel, leading to increased information loss (Section D in the supplementary material).
 These limitations become more pronounced in low-bit settings, where higher quantization errors result in performance degradation.
@@
 Acknowledgement This research was supported by the National Research Foundation (NRF) funded by the Korean government (MSIT) (No. RS-2023-00229822).
+Supplementary Material
+Appendix A Notation
+Symbol Definition
+$x$ Original activation value
+$x_{q}$ Quantized activation value
+$x_{\text{deq}}$ Dequantized activation value
+$x_{\min}$, $x_{\max}$ Minimum and maximum activation values
+$Q$ Quantization operator
+$s$ Scaling factor
+$z$ Zero-point for quantization
+$b$ Number of quantization bits
+$\min\limits_{h,w}(x)$, $\max\limits_{h,w}(x)$ Channel-wise min/max activations
+$\mathbb{R}^{C\times H\times W}$ Tensor of shape $C\times H\times W$ in $\mathbb{R}$
+[…]
+$\|\cdot\|_{2}$ L2 norm (Euclidean norm)
+$\in$ Set membership (”element of”)
+Mathematical Expressions
+Forward Quantization $x_{q}=\left\lfloor\frac{x}{s}+z\right\rceil$
+Dequantization $x_{\text{deq}}=s(x_{q}-z)$
+Scaling Factor $s=\frac{x_{\max}-x_{\min}}{2^{b}-1}$
+Zero-point $z=\left\lfloor\frac{-x_{\min}}{s}\right\rceil$
+Cosine Similarity $\frac{X\cdot Q}{\|X\|_{2}\|Q\|_{2}}$
+Relative Error $\frac{\|X-Q\|_{2}}{\|X\|_{2}}$
+Table 6: Notation for activation quantization equations, algorithm, and evaluation metrics.
+In this section, we provide an overview of the main symbols and definitions employed throughout the paper.
+It specifically includes essential elements such as the original activation ($x$), quantized activation ($x_{q}$), scaling factor ($s$), and zero-point ($z$), along with their corresponding mathematical representations.
+Appendix B Details of Experimental Setup
+We provide a detailed description of the datasets, experimental setup, and comparison methods utilized in our experiments.
+This ensures the implementation of the experiments and clearly highlights the strengths of the proposed approach over existing methods.
+B.1 Datasets
+This paper outlines the experimental settings for CIFAR-10 [9], CIFAR-100 [9], and ImageNet (ILSVRC-2012) [4] datasets.
+CIFAR-10 and CIFAR-100 contain 10 and 100 classes, respectively, with 32 × 32 resolution color images, and both have 50 K training and 10 K test images.
+ImageNet comprises 1,000 classes, with 1.2 M training and 50 K validation images, all resized to 224 × 224 for our experiments.
+We employ these widely used benchmarks for zero-shot quantization (ZSQ) to assess and compare the performance of GranQ against those of existing approaches.
+B.2 Experimental Setup
+The data generation process followed the settings of AdaDFQ [14].
+Specifically, ACGAN [12] was used for data generation with the Adam optimizer, set with a momentum of 0.9 and a learning rate of 1e-3.
+All hyper-parameters such as $\alpha_{ds}$, $\alpha_{as}$, $\lambda_{l}$, and $\lambda_{u}$ followed the default settings from AdaDFQ [14].
+For the fine-tuning process, a warmup phase was applied to ensure training stability, with 4 epochs out of 200 for CIFAR and 50 epochs out of 400 for ImageNet.
+The batch size was set to 16 for CIFAR-10 and ImageNet, and 200 for CIFAR-100.
+All experiments used Nesterov-based SGD [11] as the optimizer, with a momentum of 0.9 and a weight decay of 1e-4.
+In CIFAR experiments, the weight decay of 1e-4 was reduced by a factor of 0.1 every 100 epochs, whereas for ImageNet, decay was applied at 350 and 400 epochs (for ResNet-50 at 150, 300, and 400 epochs).
+GranQ was implemented using PyTorch [13], and visualizations were generated with matplotlib [7].
+All experiments were conducted on an NVIDIA RTX 3090 GPU.
+B.3 Comparison Methods
+GranQ is a quantization-aware training (QAT) method for zero-shot quantization (ZSQ).
+Therefore, to evaluate the proposed approach, we conducted comparative experiments using the following state-of-the-art methods:
+Algorithm 1 GranQ: Layer-Channel Quantization
+1:Activation tensors $\{x_{l}\}_{l=1}^{L}$, where $x_{l}\in\mathbb{R}^{C\times(HW)}$, quantization bit $b$
+2:Dequantized activation tensors $\{x_{\text{deq},l}\}_{l=1}^{L}$
+3:for each layer $l$ in $L$ do $\triangleright$ Process each layer separately
+4:       Compute per-channel min/max values:
+$\vec{x}_{\min}=\min_{h,w}(x_{l}),\quad\vec{x}_{\max}=\max_{h,w}(x_{l})$
+$\triangleright$ Compute along the channel axis for layer $l$
+5:       Compute scaling factors and zero-points (Vectorized):
+$\vec{s}=\frac{\vec{x}_{\max}-\vec{x}_{\min}}{2^{b}-1},\quad\vec{z}=\left\lfloor-\frac{\vec{x}_{\min}}{\vec{s}}\right\rceil$
+6:       Quantize values (Vectorized):
+$x_{q,l}=\left\lfloor\frac{x_{l}}{\vec{s}}+\vec{z}\right\rceil$
+$\triangleright$ Element-wise quantization for all channels in layer $l$
+7:       Dequantize values (Vectorized):
+$x_{\text{deq},l}=\vec{s}(x_{q,l}-\vec{z})$
+$\triangleright$ Reconstruct activation tensor for layer $l$
+8:end for
+9:Return the dequantized tensors:
+$\textbf{return}\ \{x_{\text{deq},l}\}_{l=1}^{L}$
+  • GDFQ [15]: GDFQ (Generative low-bit data-free quantization) is the first approach to train quantized models using synthetic data through knowledge distillation.
+It demonstrated that effective quantization is possible even without access to original data.
+  • AIT [3]: AIT was the first to show the negative impact of cross-entropy loss when training quantized models with distillation.
+It also provided a systematic analysis of training strategies in the ZSQ domain.
+  • AdaDFQ [14]: AdaDFQ (Adaptive data-free quantization) highlighted the generalization issues of generated data under various quantization settings and proposed a more adaptable data generation method.
+  • TexQ [2]: TexQ introduced a data generation method that preserves texture distributions, thereby enhancing the quality of generated data in ZSQ environments.
+  • RIS [1]: RIS (Robustness-guided image synthesis) addressed the low semantic quality of generated data and proposed using perturbations to improve data diversity.
+  • PLF [5]: PLF was the first to evaluate generated data before quantization, offering a method to ensure more reliable data.
+  • GenQ [10]: GenQ introduced the use of stable diffusion in ZSQ, enabling the generation of high-resolution, realistic data.
+  • AKT [6]: AKT (Advanced knowledge transfer) proposed refining feature information during distillation to improve the training of quantized models.
+  • SynQ [8]: SynQ enhanced the training of quantized models by reducing noise in generated data using low-pass filtering and aligning activation maps.
+Figure 6: Ablation study on relative error in ResNet-20 across layers.
+Layer-Q, Channel-Q, and GranQ represent layer-wise, channel-wise, and proposed quantization.
+Experiments used AdaDFQ on CIFAR-100 with 3-bit quantization.
+Appendix C Algorithm
+GranQ is a granular quantization method that processes activation values across all channels within each layer.
+Algorithm 1 describes the procedure for quantizing the activation tensor $\{x_{deq,l}\}^{L}_{l=1}$.
+The activation tensor of layer $l$ has a shape of $C\times(HW)$, and the minimum and maximum values are calculated independently for each channel.
+Based on these values, a scaling factor and zero-point are computed per channel for the quantization process.
+The resulting quantized values are dequantized back to floating-point representation prior to returning.
+Appendix D Extra Analysis
+D.1 Analysis of Activation Preservation
+To analyze the severe quantization error (activation distortion) occurring in specific layers, we conducted additional experiments, as illustrated in Figure 6.
+This figure presents the relative error for each layer of ResNet-20 and reveals that certain layers such as the 18th exhibit significantly high errors.
+These findings suggest that both layer- and channel-wise quantization fail to effectively preserve activations in specific layers.
+In contrast, GranQ achieves consistently low errors across all layers, ensuring better activation preservation.
+D.2 Analysis of Activation Magnitude
+In Figure 7, we observe that channel-wise quantization is ineffective in ZSQ settings, similar to layer-wise quantization.
+As shown in Figure 7(a), the full-precision (FP) model exhibits significant activation magnitude differences across layers, even within the same channel axis.
+This highlights the limitations of channel-wise quantization in handling quantization errors.
+Conversely, GranQ addresses this issue by accounting for each channel across all layers, thereby minimizing quantization errors.
+This is particularly important in the final layers, where maintaining the activations of the FP model is critical for accurate inference.
+Whereas (a) fails to preserve these distributions, (b) successfully maintains them.
+(a) Layer Quantization
+(b) GranQ (Ours)
+Figure 7: Visualization of activation magnitudes across layers in ResNet-20.
+(a) illustrates the activation magnitudes with layer-wise quantization, whereas (b) presents the results after applying the proposed GranQ method.
+In each subplot, the left side depicts activation magnitudes before quantization (32-bit), whereas the right side shows them after quantization (3-bit).
+The proposed GranQ effectively preserves the activation distribution, mitigating distortions caused by quantization.

### v4

@@
-GranQ: Granular Zero-Shot Quantization with Unified Layer-Channel Awareness
+GranQ: Granular Zero-Shot Quantization with Channel-Wise Activation Scaling in QAT
 Abstract
-Zero-shot quantization (ZSQ) enables neural network compression without training data, which is crucial in restricted data access environments.
-However, existing ZSQ methods suffer from significant activation loss in low-bit environments owing to their coarse-grained scaling strategy.
-To address this issue, we propose GranQ, a novel ZSQ approach that leverages layer-channel awareness to minimize the quantization error.
-Unlike conventional layer- or channel-wise quantization, GranQ dynamically adjusts quantization granularity by considering both layer- and channel-level activation distributions.
-This enables fine-grained quantization while minimizing activation distortion.
-Additionally, we introduce vectorized activation quantization, which enables efficient parallel computation and reduces computational overhead while preserving accuracy.
-GranQ achieves superior performance compared with those of state-of-the-art ZSQ methods that employ quantization-aware training.
-With these findings, we anticipate that GranQ will inspire novel research directions beyond conventional ZSQ approaches focused on data generation and model training.
-The official code will be released upon completion of the review.
+Zero-shot quantization (ZSQ) enables neural network compression without original training data, making it a promising solution for restricted data access scenarios.
+To compensate for the lack of data, recent ZSQ methods typically rely on synthetic inputs generated from the full-precision model.
+However, these synthetic inputs often lead to activation distortion, especially under low-bit settings.
+As a result, existing methods struggle to mitigate this issue due to coarse activation scaling.
+To address this issue, we propose GranQ, a novel activation quantization framework that efficiently applies per-channel scaling through vectorized computation.
+In contrast to conventional channel-wise methods, which apply vectorization only to the quantization step, GranQ improves efficiency by vectorizing the scaling operation.
+This design allows GranQ to maintain fine-grained quantization granularity with minimal computational overhead, even in low-bit environments.
+Extensive experiments under quantization-aware training (QAT) settings demonstrate that GranQ consistently outperforms state-of-the-art ZSQ methods across CIFAR and ImageNet.
+In particular, our method achieves up to 5.45% higher accuracy in the 3-bit setting on CIFAR-100 and even surpasses the full-precision baseline on CIFAR-10.
+Furthermore, GranQ achieves significant speedup in quantization latency over conventional per-channel methods, demonstrating improved efficiency.
+With these findings, we anticipate that GranQ will inspire future research beyond conventional ZSQ approaches centered on data generation and model fine-tuning.
 1 Introduction
-Figure 1: Comparison of quantization strategies.
-Existing methods apply either per-tensor scaling with layer-wise quantization or per-channel scaling.
-GranQ combines per-channel scaling with per-layer iteration for more granular quantization.
-Neural network compression techniques have been extensively studied to make large-scale deep learning (DL) models practically deployable.
+Neural network compression has been extensively studied for the practical deployment of large-scale deep learning (DL) models.
 In particular, reducing model size while minimizing performance degradation is crucial for utilizing DL on edge devices (e.g., mobile phones, embedded systems, and drones).
-Major approaches to model compression [7, 12] include quantization [24, 30, 17], pruning [28, 16, 33, 6], knowledge distillation [21, 20, 40], and neural architecture search [45, 13].
+Major approaches to model compression include quantization [1, 2, 3], pruning [4, 5, 6, 7], knowledge distillation [8, 9, 10], and neural architecture search [11, 12], as surveyed in [13, 14].
 Among these, quantization has emerged as the most actively studied technique.
-It serves as an effective compression method by reducing unnecessary representational ranges in the model while maintaining its performance.
-However, a major limitation of quantization is the need for fine-tuning or calibration to maintain the performance of the full-precision (FP) model [17].
-To address this limitation, zero-shot quantization (ZSQ) (i.e., data-free quantization) has been introduced [34, 41], which allows compressed models to maintain their performance without requiring access to the original training data.
+It serves as an effective compression method by reducing unnecessary representational ranges in the model.
+However, it often requires fine-tuning or calibration to match full-precision (FP) model performance [3].
+To address this, zero-shot quantization (ZSQ), also known as data-free quantization, has been proposed to compress models without the original training data.
 Since the introduction of ZeroQ [15], studies on ZSQ have advanced in two main directions.
@@
 We categorize existing ZSQ methods based on these directions, as summarized in Table 1.
+First, data generation studies focus on generating high-quality data to effectively train the Q model [16, 17, 18].
+Meanwhile, calibration (PTQ) studies aim to minimize the quantization error by calibrating the Q model without additional training [19, 20].
+Finally, fine-tuning (QAT) studies focus on transferring key information from the FP model to the Q model through knowledge distillation [21, 22, 23, 24].
+However, despite extensive studies, severe performance degradation in low-bit quantization remains unresolved.
+To address this issue, we performed an in-depth analysis of the ZSQ process, focusing on why low-bitwidth settings still suffer from performance loss even after QAT fine-tuning.
+Our findings reveal that quantization errors mainly stem from the loss of activation values instead of data quality or training methods.
+Notably, we found that layer-wise (per-tensor) quantization is no longer suitable for preserving activations in ZSQ, as it leads to coarse and inaccurate representations.
+Based on this analysis, we introduce GranQ, a novel ZSQ approach that enables channel-wise activation quantization by adapting the scale to each channel of every input activation.
+This allows us to overcome the limitations of conventional activation quantization.
+This dynamic adjustment minimizes activation loss and preserves the original activation values by reducing quantization errors, as shown in Figure 1.
+The proposed method effectively handles activation loss in low-bit quantization and achieves state-of-the-art (SOTA) performance in QAT settings on the CIFAR and ImageNet datasets.
+Furthermore, We apply vectorization to the scaling operation, which is not applied in conventional channel-wise quantization, reducing latency while enabling fine-grained activation quantization.
+Our contributions can be summarized as follows:
+(a) Layer-wise Quantization
+(b) GranQ (Channel-wise, Ours)
+Figure 1: Comparison between (a) layer-wise quantization and (b) GranQ on the CIFAR-10.
+Each subfigure visualizes the 32-bit FP (left) and 3-bit quantized (right) activations of the first ResNet-20 layer.
+GranQ better preserves the original activation with minimal distortion.
+  • We identify critical limitations of layer-wise activation quantization in low-bit ZSQ.
+Our findings reveal that conventional activation quantization methods relying on layer-wise granularity suffer from significant activation loss.
+This limitation becomes more severe in ZSQ settings with synthetic data.
+  • We propose GranQ, a novel method that supports granular quantization and maintains computational efficiency.
+We introduce a fine-grained activation quantization mechanism that considers every channel to reduce activation loss.
+Although per-channel activation quantization is known to improve precision, it has not been widely adopted due to its high computational cost.
+We address this by applying vectorization to the scaling operation, enabling efficient per-channel quantization with minimal overhead.
+To the best of our knowledge, our approach is the first to address the ZSQ problem.
+  • We achieve SOTA performance over existing ZSQ methods through extensive evaluation.
+Specifically, on the CIFAR-100 dataset, in the 3-bit quantization setting, GranQ achieves an accuracy of 62.73%, improving by 5.45% over the latest method on ResNet-20.
+Furthermore, on the CIFAR-10 dataset in the 5-bit quantization setting, GranQ achieves an accuracy of 94.06%, slightly exceeding the FP model performance by 0.17%.
+2 Related Work
+2.1 Quantization
+Quantization reduces the representational range of deep neural networks (DNNs) and minimizes memory usage.
+It is typically categorized into post-training quantization (PTQ) and quantization-aware training (QAT) [3, 25].
+PTQ applies quantization after training without further updates, which makes it efficient and easy to implement.
+However, it is sensitive to scaling errors and often relies on calibration to improve accuracy.
+In contrast, QAT incorporates quantization during training and optimizes quantized activations using the straight-through estimator [3, 26].
+Both methods require data—PTQ for calibration and QAT for fine-tuning.
+In practice, access to training data is often limited.
+To address this, zero-shot quantization (ZSQ) has been proposed to perform quantization without using original data [16].
 Data Generation
@@
 Data generation methods typically support both PTQ and QAT, whereas violet-marked methods utilize fine-tuning for data generation and support only QAT.
-First, data generation studies focus on generating high-quality data to effectively train the Q model.
-A common approach uses batch normalization statistics from the FP model [41].
-Recent studies focused on generating data adaptable to various bit-widths [36] and using diverse generative models such as stable diffusion guided by text prompts [32].
-Meanwhile, calibration (PTQ) studies aim to minimize the quantization error by calibrating the Q model without additional training [18, 2].
-Finally, fine-tuning (QAT) studies focus on transferring key information from the FP model to the Q model through knowledge distillation [10, 14, 23, 26].
-However, despite extensive studies, severe performance degradation in low-bit quantization remains unresolved.
-To address this issue, we performed an in-depth analysis of the ZSQ process, focusing on why low-bitwidth settings still suffer from performance loss even after QAT fine-tuning.
-Our findings reveal that quantization errors mainly stem from the loss of activation values instead of data quality or training methods.
-Notably, we discovered that layer-wise (per-tensor) quantization fails to accurately preserve activations in ZSQ, leading to coarse and inaccurate representations.
-Based on this analysis, we introduce GranQ, a novel ZSQ approach.
-GranQ independently handles large activations in specific layers and channels by narrowing the range between the maximum and minimum values within each channel (Figure 1).
-This dynamic adjustment minimizes activation loss and preserves the original activation values by reducing quantization errors, as shown in Figure 2.
-The proposed method effectively handles activation loss in low-bit quantization and achieves state-of-the-art (SOTA) performance in QAT settings on the CIFAR and ImageNet datasets.
-Furthermore, we implement vectorized operations for granular quantization with minimal computational overhead, enabling efficient quantization on GPUs.
-Our contributions can be summarized as follows:
-  • We perform a comprehensive analysis of quantization errors in low-bit ZSQ settings.
-Our findings reveal that conventional activation quantization methods (e.g., layer- and channel-wise) lead to significant activation loss.
-This issue stems from coarse-grained activation quantization, which fails to preserve fine-grained details.
-  • We propose GranQ, a novel method that supports granular quantization and maintains computational efficiency.
-We introduce a fine-grained quantization mechanism, which considers every channel in each layer, to reduce activation loss.
-Additionally, we mitigate computational overhead using vectorized activation quantization.
-To the best of our knowledge, our approach is the first to address the ZSQ problem.
-  • We achieve SOTA performance over existing ZSQ methods through extensive evaluation.
-Specifically, on the CIFAR-100 dataset, in the 3-bit quantization setting, GranQ achieves an accuracy of 62.73%, improving by 5.45% over the latest method on ResNet-20.
-Furthermore, on the CIFAR-10 dataset in the 5-bit quantization setting, GranQ achieves an accuracy of 94.06%, slightly exceeding the FP model performance by 0.17%.
-(a) Layer Quantization
-(b) GranQ (Ours)
-Figure 2: Comparison between (a) layer-wise quantization and (b) the proposed GranQ method on the CIFAR-10 dataset from AdaDFQ[36].
-We visualize the activation of the first layer in ResNet-20 for the first image of the first batch.
-$w$ and $a$ denote weight and activation, respectively.
-Each sub-figure presents the activation graph for 32-bit FP (left) and 3-bit quantization (right).
-Although layer-wise quantization significantly distorts the original activation after quantization, GranQ effectively preserves it with minimal distortion.
-2 Related Work
-2.1 Quantization
-Quantization can reduce the representational range of pre-trained deep neural networks (DNNs) and minimize memory usage.
-Quantization methods are categorized into PTQ and QAT based on the timing of the range reduction.
-PTQ quantizes fully trained DNNs without any further training [17, 3], offering low computational cost and simple implementation.
-However, because PTQ directly applies quantization to the FP model, it struggles to properly adjust the range during quantization.
-To overcome this, calibration is used to select optimal quantization ranges by determining scaling factors.
-In contrast, QAT fine-tunes the Q model with quantization operations integrated into the training process, enabling direct optimization of activations [17].
-In particular, it employs the straight-through estimator [42] to estimate gradients.
-However, both PTQ and QAT have the limitation of requiring access to data.
-PTQ needs data for calibration, whereas QAT requires it for fine-tuning.
-In practice, obtaining training data is often challenging.
-To address this problem, ZSQ has been proposed, which performs quantization without using original data [41].
 2.2 Zero-shot Quantization
@@
 Recent studies have focused on generating high-quality synthetic data [31, 18] and developing more effective methods to utilize them [21, 23, 24].
-However, these approaches are still limited in preserving activations in low-bit settings.
-Therefore, a new quantization method is required to effectively minimize activation loss.
-In this study, we propose a novel approach to address this challenge.
+However, current ZSQ methods heavily rely on augmented synthetic inputs.
+This causes large variation in activation scales across channels.
+In practice, such input-dependent channel-wise quantization is rarely adopted due to its high computational cost.
+Therefore, recent studies have applied channel-wise activation quantization either in a limited conditions within QAT settings [35], or exclusively in PTQ scenarios with access to a small amount of data [36, 37].
+This highlights the need for efficient channel-wise activation quantization for QAT in ZSQ.
 3 Preliminaries and Problem Definition
-In this section, we provide a new definition of the existing ZSQ problem and introduce the preliminaries. The detailed mathematical notations are listed in Table 6 in the supplementary material.
+In this section, we provide a new definition of the existing ZSQ problem and introduce the preliminaries. The detailed mathematical notations are listed in Table 6 in the appendix.
 3.1 Activation Quantization
-Activation quantization reduces the precision of intermediate activation values in neural networks by converting them into low-bitwidth representations [8].
-This technique is often used alongside weight quantization to further compress the model.
-Among various activation quantization methods, linear quantization is the most widely used.
-It uses a normalized scaling factor to adjust the activation values, and subsequently converts them into integers within a fixed bit range [24, 15].
-The overall process typically involves three stages: forward quantization, scaling parameter computation, and dequantization [17].
-3.1.1 Forward Quantization
-The forward quantization process involves mapping continuous floating-point values to an integer space. This process utilizes the quantization operator $Q$, and the equation is defined as follows:
+Activation quantization reduces the precision of intermediate activation values by converting them into low-bitwidth integers [38].
+It is commonly used with weight quantization to compress models, and the linear quantization scheme is most widely adopted.
+This process typically consists of quantization, scaling parameter computation, and dequantization [1, 3, 39].
+The quantization operator $\mathcal{Q}$ maps a floating-point value $x$ into an integer $x_{q}$ using a scaling factor $s$ and zero-point $z$:
 $x_{q}=\mathcal{Q}(x,s,z)=\left\lfloor\frac{x}{s}+z\right\rceil$ (1)
-[…]
-3.1.2 Scaling Parameter Computation
-In activation quantization, setting the scaling factor $s$ is crucial for mapping various activation distributions to a fixed integer range.
-In this study, we employ the asymmetric quantization method, which is the most commonly used for activation quantization [29].
-We calculate the scaling factor $s$ and the zero-point $z$ as follows:
+Here, $s$ normalizes the activation range, and $z$ shifts the scaled value to account for asymmetric quantization. These parameters are computed as:
 $s=\frac{x_{\max}-x_{\min}}{2^{b}-1},\quad z=\left\lfloor\frac{-x_{\min}}{s}\right\rceil$ (2)
-Here, $x_{\max}$, $x_{\min}$ and $b$ represent the maximum and minimum values of the activation, and the number of quantization bits, respectively.
-The zero-point $z$ represents the offset used to adjust activation values.
-$x_{\max}$ and $x_{\min}$ determine the range of the activation distribution, which directly affects the computation of scaling factor $s$.
-Specifically, when $x_{\max}-x_{\min}$ is large, the scaling factor $s$ increases.
-This leads to a larger quantization step size, making it harder to preserve fine-grained activations.
-Conversely, when $x_{\max}-x_{\min}$ is small, the scaling factor $s$ also decreases.
-This allows for finer value distinctions and enhances the quality of representation.
-3.1.3 Dequantization
-The dequantization process converts quantized integer values back into their real-valued representation, as shown in the following equation:
+where $x_{\min}$ and $x_{\max}$ denote the range of activation values, and $b$ is the bit-width.
+A large range yields a large $s$, increasing quantization error; a smaller range allows finer resolution.
+During inference, the dequantization step restores quantized values to real numbers:
 $x_{\text{deq}}=s(x_{q}-z)$ (3)
-In this process, the zero-point $z$ is subtracted to reverse the forward quantization, and the result is multiplied by the scaling factor $s$ to obtain the final value $x_{deq}$.
-The dequantization process restores quantized activation values to their original real-valued range.
-It is essential during model inference for converting final predictions into real values.
-This process is essential for converting the final prediction into real values during model inference.
-Figure 3: Visualization of activation magnitudes for the first image in the first batch using ResNet-20 on CIFAR-10.
-The left column illustrates activations from the FP model, whereas the right column presents activations after layer-wise quantization.
-Significant distortions are observed in certain layers, particularly in the final layer.
+This enables recovery of activation information needed for final predictions.
 3.2 Problem Definition
 Existing activation quantization methods are designed based on layer-wise quantization to minimize computational cost.
-However, fixed scaling factors struggle to handle varying activations.
+However, a fixed scaling factor struggles to handle varying activations.
 This issue becomes more severe under low-bitwidth settings in ZSQ environments.
 The following outlines the primary challenges faced in ZSQ under low-bitwidth settings.
-(Problem 1) Coarse-grained activation representation
-In conventional activation quantization approaches, the scaling factor $s$ is fixed per layer or channel, mapping all activations to the same range.
-However, in low-bitwidth settings, this approach struggles to capture fine-grained activation variations from generated data.
+(Problem 1) Coarse Activation Quantization by Single-Range Scaling
+Conventional activation quantization uses a single scaling range per layer for efficiency, which was acceptable at higher bitwidths or when real data is available.
+However, this design becomes problematic in ZSQ, where range adjustment relies on synthetic data.
+These synthetic distributions are often biased or irregular, making precise quantization more difficult.
 (Problem 2) Severe Quantization Error in Certain Layers
-During quantization, certain layers induce severe quantization errors, leading to significant activation distortions.
+During activation quantization, certain layers induce severe quantization errors, leading to significant activation distortions.
 These distortions substantially differ from the original distribution and can directly impact the performance of the model.
@@
 We observe the causes of activation distortion during quantization and analyze how to mitigate them. Based on this, we introduce the proposed method known as GranQ.
+Figure 2: Activation magnitudes of ResNet-20 on CIFAR-10 using 32-bit FP (left) and 3-bit layer-wise quantization (right). Severe distortions appear, especially in the final layer.
 4.1 Observation
-ResNet-20 Layer-Q Channel-Q GranQ
-$Avg.\frac{X\cdot Q}{\|X\|_{2}\|Q\|_{2}}(\uparrow)$
-$Avg.\frac{\|X-Q\|_{2}}{\|X\|_{2}}(\downarrow)$
-0.5111
-0.3129
-0.5602
-0.2528
-0.6835
-0.1063
-Table 2: Comparative analysis of averaged cosine similarity and relative error over all layers and batches.
-Higher cosine similarity ($\uparrow$) indicates better retention of activation information, whereas lower relative error ($\downarrow$) signifies reduced quantization distortion.
-Layer-Q and Channel-Q denote layer- and channel-wise quantization, respectively, whereas GranQ integrates both for improved quantization.
 Layer-wise quantization effectively captures the activation distribution of each layer.
@@
 We analyzed this phenomenon occurring across all layers and provided a summary in Table 2.
-In Table 2, we present the cosine similarity $\frac{X\cdot Q}{|X|_{2}|Q|_{2}}$, along with the relative error $\frac{|X-Q|_{2}}{|X|_{2}}$.
-From Table 2, we identify that both layer- and channel-wise quantization fail to preserve the original activation effectively.
-This indicates that layer- and channel-wise quantization, which independently process activations at the layer or channel level, are not suitable for ZSQ.
-In contrast, GranQ, which considers both layers and channels simultaneously, achieves 1.34$\times$ higher similarity and 2.94$\times$ lower relative error compared with those of layer-wise quantization, demonstrating its effectiveness in preserving activations.
+It presents the cosine similarity $\frac{X\cdot Q}{|X|_{2}|Q|_{2}}$, along with the relative error $\frac{|X-Q|_{2}}{|X|_{2}}$.
+From the results, we observe that layer-wise quantization fails to preserve activation information effectively.
+This reveals that using a single scaling range for activation quantization has a critical limitation, particularly in ZSQ.
+In contrast, GranQ, which applies scaling per channel, achieves 1.34$\times$ higher similarity and 2.94$\times$ lower relative error than layer-wise quantization, demonstrating its effectiveness in preserving activation information.
 Furthermore, Figure 2 provides a visual representation of the effect of quantization on activation values across layers.
-In layer-quantizaiton, although the overall activations generally reflect the FP activations to some extent, certain layers exhibit fine activation values being mapped to discrete values, leading to coarse representations.
+In layer-wise quantizaiton, although the overall activations generally reflect the FP activations to some extent, certain layers exhibit fine activation values being mapped to discrete values, leading to coarse representations.
 These coarse representations have been identified as a major factor leading to a decline in the overall expressiveness and inference performance of the model.
-Further analysis of Figure 3 is provided in the supplementary material, in Figures 6 and 7.
-4.2 Granular Quantization
-Based on our analysis, we propose GranQ (granular layer-channel quantization), a fine-grained ZSQ approach designed to minimize computational overhead, as shown in Figure 4.
-GranQ combines layer and channel information using layer-channel scaling and improves efficiency through vectorized quantization.
-4.2.1 Layer-Channel Scaling
-We propose layer-channel scaling, which enables GranQ to assign individual scaling factors to each channel, enhancing quantization precision.
-This approach reduces the impact of outlier activations in specific channels on the quantization process.
-The layer-channel scaling is defined as follows:
-$\vec{x}_{\min}$ $=\min_{h,w}(x_{l}),\quad\vec{x}_{\max}=\max_{h,w}(x_{l})$ (4)
-where $\text{where }x_{l}\in\mathbb{R}^{C\times H\times W}$
+Further analysis of Figure 2 is provided in Figure 5 and in Appendix (Figure 7).
+Activaiton Quantization Layer-wise Channel-wise (GranQ)
+$Avg.\frac{X\cdot Q}{\|X\|_{2}\|Q\|_{2}}(\uparrow)$
+$Avg.\frac{\|X-Q\|_{2}}{\|X\|_{2}}(\downarrow)$
+0.5111
+0.3129
+0.6835
+0.1063
+Table 2: Average cosine similarity and relative error for 3-bit activation quantization on ResNet-20 (CIFAR-100).
+GranQ (channel-wise) outperforms layer-wise quantization by achieving less distortion in activation representation, as indicated by higher similarity and lower error.
+4.2 Methodology
+Based on our analysis, we propose GranQ, a fine-grained ZSQ method designed to reduce computational overhead (Figure 3).
+GranQ applies channel-wise scaling through activation reshaping and decomposition, and achieves fast parallel processing via vectorization of scaling parameters, which enables efficient computation of both scaling and quantization.
+4.2.1 Activation Reshaping and Decomposition
+We propose activation reshaping and decomposition, which reshapes each activation map from a three-dimensional tensor of shape $(C\times H\times W)$ into a two-dimensional matrix of shape $(C\times HW)$.
+This transformation enables decomposition of the activation map along the channel axis, allowing each channel to be processed independently.
+It also facilitates vectorized computation of per-channel statistics (e.g., min and max) required for scaling factor calculation during quantization.
+The activation reshaping and decomposition is defined as follows:
+$\begin{gathered}A_{l}=\left[A_{l}(1,:,:),\,A_{l}(2,:,:),\,\dots,\,A_{l}(C,:,:)\right]\in\mathbb{R}^{C\times H\times W}\\ \vec{A}_{\min}=\left[\min_{(h,w)}A_{l}(c,h,w)\right]_{c=1}^{C}\in\mathbb{R}^{C}\\ \vec{A}_{\max}=\left[\max_{(h,w)}A_{l}(c,h,w)\right]_{c=1}^{C}\in\mathbb{R}^{C}\end{gathered}$ (4)
 […]
+4.2.2 Vectorization of Scaling Parameters
+In the vectorization of scaling parameters stage, multiple scaling ranges are stored in a vectorized form for each channel.
+This allows dynamic adaptation to the diverse activation distributions induced by synthetic inputs.
+Thus, GranQ contrasts with traditional per-channel quantization, which uses fixed scaling parameters for each channel regardless of input variation.
+As shown in Equation 5, the scaling factors $\vec{s}$ and zero-points $\vec{z}$ are computed in vectorized form.
+The activation values are then quantized and dequantized according to Equation 6, with all operations applied in parallel across channels.
+$\vec{s}=\frac{\vec{A}_{\max}-\vec{A}_{\min}}{2^{b}-1},\quad\vec{z}=\left\lfloor-\frac{\vec{A}_{\min}}{\vec{s}}\right\rceil$ (5)
+$A_{q,l}=\left\lfloor\frac{A}{\vec{s}}+\vec{z}\right\rceil,A_{\text{deq},l}=\vec{s}(A_{q,l}-\vec{z})$ (6)
+This process allows efficient and adaptive quantization of activation tensors exhibiting high variation across inputs, leading to improved computational efficiency (Table 7 in appendix).
+[…]
+5 Experiments
+In this section, we thoroughly evaluate the effectiveness of GranQ. Experiments are conducted on diverse benchmark datasets, with the performance compared with those of existing ZSQ methods.
+5.1 Experimental Setup and Details
+The experiments were conducted using widely adopted ZSQ evaluation datasets, including CIFAR-10, CIFAR-100 [40], and ImageNet (ILSVRC 2012) [41] validation datasets.
+For the CIFAR datasets, ResNet-20 [42] was used as the quantization model, whereas ResNet-18 [42], ResNet-50 [42], and MobileNetV2 [43] were employed for ImageNet.
+All experiments were conducted using the SGD optimizer [44] with a momentum of 0.9 and weight decay of 1e-4.
+The CIFAR-10 and CIFAR-100 experiments were each conducted for 200 epochs, with batch sizes of 16 and 200, respectively.
+For ImageNet, we trained for 400 epochs with a batch size of 16.
+The initial learning rate was set to 1e-4 for CIFAR-10 and CIFAR-100, and 1e-5 for ImageNet, with multi-step learning rate decay applied.
+The decay steps were set to 100, 200, and 300 epochs for CIFAR, and at 350 and 400 epochs for ImageNet, with a decay rate of 0.1.
+We compared our method with existing ZSQ methods [16, 21, 17, 33, 34, 22, 18, 23, 24].
+For data generation, we followed the AdaDFQ [17] approach based on ACGAN [45].
+Further implementation details are provided in the appendix, Section A.3.2 and A.3.3.
+In our ablation study, layer-wise quantization was applied to all layers containing activation functions, while channel-wise quantization was performed per channel at the batch level.
 Dataset
@@
 (AAAI 24)
-PLF
-(CVPR 24)
 GenQ
@@
 $\dagger$ indicates our re-implementation with the official code.
-[…]
-4.2.2 Vectorized Quantization
-In the layer-channel scaling stage, the activation tensor of each layer is vectorized along the channel dimension.
-In the subsequent vectorized quantization stage, both the scaling factor $\vec{s}$ and zero-point $\vec{z}$ are also vectorized and computed as shown in Equation 5.
-The activation values go through quantization and dequantization (Equation 6), with all operations running in parallel across channels.
-During computation, scaling factor $s$ and zero-point $z$ of each channel are broadcasted to the activation values and applied element-wise.
-$\vec{s}=\frac{\vec{x}_{\max}-\vec{x}_{\min}}{2^{b}-1},\quad\vec{z}=\left\lfloor-\frac{\vec{x}_{\min}}{\vec{s}}\right\rceil$ (5)
-$x_{q,l}=\left\lfloor\frac{x}{\vec{s}}+\vec{z}\right\rceil,x_{\text{deq},l}=\vec{s}(x_{q,l}-\vec{z})$ (6)
-This process can be applied to every layer in the network, enabling the vectorized quantization of all activation tensors.
-A detailed computational process is provided in Algorithm 1, available in the supplementary material.
-5 Experiments
-In this section, we thoroughly evaluate the effectiveness of GranQ. Experiments are conducted on diverse benchmark datasets, with the performance compared with those of existing ZSQ methods.
-5.1 Experimental Setup and Details
-The experiments were conducted using widely adopted ZSQ evaluation datasets, including CIFAR-10, CIFAR-100 [27], and ImageNet (ILSVRC 2012) [11] validation datasets.
-For the CIFAR datasets, ResNet-20 [19] was used as the quantization model, whereas ResNet-18 [19], ResNet-50 [19], and MobileNetV2 [38] were employed for ImageNet.
-All experiments were conducted using the SGD optimizer [37] with a momentum of 0.9 and weight decay of 1e-4.
-The CIFAR-10 and CIFAR-100 experiments were each conducted for 200 epochs, with batch sizes of 16 and 200, respectively.
-For ImageNet, we trained for 400 epochs with a batch size of 16.
-The initial learning rate was set to 1e-4 for CIFAR-10 and CIFAR-100, and 1e-5 for ImageNet, with multi-step learning rate decay applied.
-The decay steps were set to 100, 200, and 300 epochs for CIFAR, and at 350 and 400 epochs for ImageNet, with a decay rate of 0.1.
-We compared our method with existing ZSQ methods [41, 10, 36, 5, 1, 14, 32, 23, 26].
-For data generation, we followed the AdaDFQ [36] approach based on ACGAN [35].
-Further implementation details are provided in the supplementary material, Section B.
-In our ablation study, layer-wise quantization was applied to all layers containing activation functions, while channel-wise quantization was performed per channel at the batch level.
 CIFAR-100 ResNet-20 (70.33%)
@@
 Baseline +GranQ Baseline +GranQ Baseline +GranQ Baseline +GranQ
-3w3a 47.61 59.04+11.43 45.70 60.42+14.72 52.74 62.73+9.99 54.68 62.01+7.33
-4w4a 63.39 66.97+3.58 65.80 68.08+2.28 66.81 68.79+1.98 66.94 68.77+1.83
-5w5a 66.12 68.96+2.84 69.26 70.14+0.88 69.93 70.05+0.12 69.75 70.21+0.46
 […]
-Layer-Q Channel-Q
-Layer-Channel
-Layer-Channel
-(Vectorization, GranQ)
-Accuracy (%)
-49.98 (±$\pm 0.24$)
-51.26 (±$\pm 0.28$)
-62.68 (±$\pm 0.11$)
-62.73 (±$\pm 0.17$)
-Table 5: Comparative analysis of four quantization approaches on ResNet-20 accuracy.
-The experiments were conducted using the AdaDFQ method on CIFAR-100 with 3-bit quantization.
-Results are averaged over three repetitions with standard deviation.
+[…]
+[…]
+Table 4: Accuracy of existing SOTA methods with the integration of GranQ. GDFQ and AdaDFQ focus on data generation, whereas Qimera+AIT and AdaDFQ+AKT primarily enhance quantized model training.
+Figure 4: Latency of ResNet-20 quantization across batch sizes on CIFAR-100 with 3-bit setting.
+Figure 5: Relative quantization error across layers in ResNet-20 with 3-bit quantization on CIFAR-100.
 5.2 Performance Evaluation
@@
 However, GenQ uses a pre-trained diffusion model [46] and prompts during data generation, leading to slower speeds compared with those of typical GAN-based methods [16].
-In contrast, GranQ maintains the speed advantage of GAN-based generation while achieving high performance through fine-grained adjustments.
+In contrast, GranQ improves the quantization mechanism without relying heavily on diffusion-based data generation, as GenQ does.
+As our method is adaptable to existing QAT frameworks, it can potentially achieve even higher performance when combined with GenQ’s data generation.
 5.3 Ablation Study
@@
 These findings demonstrate that GranQ is not only effective in data synthesis-based methods but also enhances performance during the model training process.
-Figure 5: Ablation study on the latency of the quantization process in ResNet-20 across different batch sizes.
-The experiments were conducted using the AdaDFQ method on CIFAR-100 with a 3-bit quantization setting.
 5.3.2 Efficiency Evaluation
-In general, performing precise quantization that considers both layers and channels leads to an exponential increase in computational cost.
-However, GranQ is designed to minimize this overhead while maintaining high accuracy by leveraging vectorized operations.
-To validate this, we measured the latency across various batch sizes (16, 32, 64, 128, and 200).
-Figure 5 compares four methods.
-The first two methods are conventional quantization techniques: layer- and channel-wise.
-The third is scalar-based iterative, which considers all channels in each layer but performs multiple iterations.
-The fourth method is GranQ, which integrates both layer and channel information while employing vectorized quantization.
-Although scalar-based iterative quantization performs the same operations as GranQ, it relies on repeated scalar computations, as illustrated in Equations 1, 2, and 3.
-In contrast, GranQ enhances efficiency through parallelized computations (Algorithm 1 in supplementary material).
-A key observation from Figure 5 is that GranQ induces only a slight increase in latency compared with layer- and channel-wise quantization.
-Moreover, as summarized in Table 5, GranQ achieves significantly higher accuracy than traditional quantization methods.
-Even when compared with the less efficient iteration-based layer-channel quantization, GranQ maintains nearly identical accuracy.
-Additionally, as summarized in Table 2, GranQ substantially reduces quantization error and effectively preserves activations.
-These findings confirm that GranQ achieves fine-grained quantization while benefiting from parallel computation.
+Method
+Quantization
+Application
+Scaling Factor
+Computation
+Accuracy (%)
+Quantization
+Latency (ms)
+Layer-wise 49.98 ($\pm$ 0.24) 0.227
+Channel-wise (Conventional scalar-loop) 62.68 ($\pm$ 0.11) 103.671
+Channel-wise (Vectorized, GranQ) 62.73 ($\pm$ 0.17) 0.236
+Table 5: Ablation study of vectorization in scaling computation and quantization.
+GranQ applies vectorization to both stages, achieving high accuracy with low latency.
+Latency is measured with batch size 16.
+Experiments use AdaDFQ on ResNet-20 (CIFAR-100, 3-bit).
+While fine-grained activation quantization often incurs high computational cost due to per-channel scaling, GranQ maintains efficiency by vectorizing the scaling computation.
+By reshaping activation tensors, our method eliminates the overhead of iterative channel-wise operations.
+To assess this in practice, we measured quantization latency across various batch sizes (16, 32, 64, 128, and 200).
+Figure 5 compares three methods: layer-wise quantization, channel-wise quantization with scalar-loop scaling (Algorithm 2 in appendix), and our method (GranQ), which efficiently parallelizes scaling across channels (Algorithm 1 in appendix).
+The two channel-wise methods perform per-channel quantization.
+The conventional method computes scaling factors sequentially, whereas GranQ vectorizes the scaling computation across channels for improved efficiency.
+A key observation from Figure 5 is that GranQ incurs only a slight increase in latency compared to conventional layer-wise quantization.
+With only a minor increase in latency, GranQ achieves higher accuracy than the layer-wise baseline, as shown in Table 5.
+Compared to scalar-loop-based channel quantization, which implements the same computation sequentially, GranQ achieves nearly identical accuracy.
+Furthermore, Table 2 shows that GranQ substantially reduces quantization error and more effectively preserves activation information.
+These results demonstrate that GranQ enables fine-grained quantization with significantly improved efficiency.
+Additional analysis and experimental results on inference efficiency are provided in Table 7 of the appendix.
 6 Discussion
 6.1 Why GranQ is Effective?
-The most common activation quantization approaches are known for their low computational cost and fast processing speeds.
+The most common approaches to activation quantization (i.e., layer-wise quantization) are known for their low computational cost and fast processing speed.
 However, they often fail to accurately capture activation distributions, resulting in substantial quantization errors (Figure 2).
 In layer-wise quantization, representing the entire activation range with a single scaling factor makes it difficult to reflect fine-grained distribution changes.
-Similarly, channel-wise quantization encounters challenges owing to significant range differences in activations across layers for the same channel, leading to increased information loss (Section D in the supplementary material).
-These limitations become more pronounced in low-bit settings, where higher quantization errors result in performance degradation.
-To overcome these issues, GranQ introduces the layer-channel scaling method, which combines layer- and channel-wise quantization.
-Unlike conventional methods, GranQ considers the activation distribution of each channel individually while also incorporating layer information, enabling more precise quantization.
-Furthermore, although conventional quantization relies on scalar operations with a single scaling factor, GranQ utilizes multiple scaling factors and performs vectorized computations.
-This approach leverages GPU resources efficiently, minimizing computational overhead while achieving effective quantization.
-Experimental results demonstrate that GranQ outperforms traditional layer- and channel-wise quantization methods, achieving higher cosine similarity and lower relative error (Table 2).
-This demonstrates that GranQ accurately preserves the original activation distribution.
-Additionally, GranQ maintains computational efficiency in terms of latency and effectively balances the trade-off between computational cost and performance (Figure 5 and Table 5).
-Most notably, GranQ achieves SOTA performance compared with those of existing layer-wise quantization methods (Table 3), highlighting its efficiency and superiority over scalar iteration approaches.
+These limitations become more severe in low-bit settings, especially when synthetic data exhibit distributional shifts.
+To overcome these issues, GranQ computes and stores per-channel scaling factors for each input in a unified vector form, enabling parallel processing across all channels.
+Unlike conventional methods, it performs quantization based on each input’s channel-wise distribution, resulting in more precise and efficient computation.
 6.2 Limitation and Future-work
-As GranQ aims to precisely preserve activations during quantization, it partially depends on generated data quality.
-Consequently, studies on data generation remain highly important.
-The proposed method has the potential to achieve even better performance when applied to high-quality data generation strategies.
-In future work, we aim to improve the quantization performance by focusing on both the accurate generation of original activations and their effective preservation.
+In ZSQ settings, synthetic data often exhibit broader and more distorted activation distributions compared to original data (Figure 6 in appendix).
+GranQ addresses this challenge by applying dynamic quantization for each channel based on its activation distribution, making it particularly effective in QAT scenarios.
+However, GranQ is less suitable for PTQ, which relies on static activation quantization.
+As future work, we plan to develop quantization methods that preserve the fine-grained scaling of GranQ while enhancing adaptability to PTQ settings.
 7 Conclusion
-We present a comprehensive analysis of activation quantization errors in ZSQ and propose GranQ, a novel quantization method to effectively reduce these errors.
-Conventional layer- and channel-wise quantization methods rely on a single scaling factor, which often results in substantial activation loss.
-In contrast, GranQ employs a fine-grained quantization approach that simultaneously considers both layer and channel information, thereby reducing activation loss and enabling more precise quantization.
-Moreover, GranQ successfully maintains computational efficiency through vectorized quantization while performing detailed and fine-grained quantization.
+Layer-wise and channel-wise quantization are both commonly used in weight quantization, as weights remain fixed after training.
+In contrast, channel-wise scaling in activation quantization is more challenging, as activation values vary with each input, making it difficult to apply fine-grained scaling.
+These limitations become more pronounced in QAT settings, where large amounts of data are used.
+Under these conditions, quantization is strongly affected by activation distributions.
+This issue is especially critical in ZSQ, where training data are synthesized using internal statistics from the FP model.
+Synthetic data generated in this way often produce distorted activation distributions compared to original data, resulting in increased quantization error during scaling.
+To address these issues, we propose GranQ, a dynamic channel-wise quantization method that adapts scaling factors to each input’s activation distribution.
+This enables more precise quantization even under distorted distributions, as often observed in ZSQ.
+In addition, GranQ improves computational efficiency by applying per-input, per-channel scaling through vectorized operations, which is rarely adopted in activation quantization due to its implementation complexity.
 Through extensive experiments, we demonstrate that GranQ achieves superior performance while preserving computational efficiency.
@@
 Our study highlights the importance of activation preservation in ZSQ and proposes GranQ as an insightful direction to address this challenge.
-Acknowledgement This research was supported by the National Research Foundation (NRF) funded by the Korean government (MSIT) (No. RS-2023-00229822).
-Supplementary Material
-Appendix A Notation
+Appendix A Appendix
+A.1 Notation
+We provide an overview of the main symbols and definitions employed throughout the paper.
+It specifically includes essential elements such as the original activation ($A$), quantized activation ($A_{q}$), scaling factor ($s$), and zero-point ($z$), along with their corresponding mathematical representations.
 Symbol Definition
-$x$ Original activation value
-$x_{q}$ Quantized activation value
-$x_{\text{deq}}$ Dequantized activation value
-$x_{\min}$, $x_{\max}$ Minimum and maximum activation values
+$A$ Original activation value
+$A_{q}$ Quantized activation value
+$A_{\text{deq}}$ Dequantized activation value
+$A_{\min}$, $A_{\max}$ Minimum and maximum activation values
 $Q$ Quantization operator
@@
 $b$ Number of quantization bits
-$\min\limits_{h,w}(x)$, $\max\limits_{h,w}(x)$ Channel-wise min/max activations
+$\min\limits_{h,w}(A)$, $\max\limits_{h,w}(A)$ Channel-wise min/max activations
 $\mathbb{R}^{C\times H\times W}$ Tensor of shape $C\times H\times W$ in $\mathbb{R}$
@@
 $\|\cdot\|_{2}$ L2 norm (Euclidean norm)
-$\in$ Set membership (”element of”)
+$\in$ Set membership ("element of")
 Mathematical Expressions
@@
 Table 6: Notation for activation quantization equations, algorithm, and evaluation metrics.
-In this section, we provide an overview of the main symbols and definitions employed throughout the paper.
-It specifically includes essential elements such as the original activation ($x$), quantized activation ($x_{q}$), scaling factor ($s$), and zero-point ($z$), along with their corresponding mathematical representations.
-Appendix B Details of Experimental Setup
-We provide a detailed description of the datasets, experimental setup, and comparison methods utilized in our experiments.
-This ensures the implementation of the experiments and clearly highlights the strengths of the proposed approach over existing methods.
-B.1 Datasets
+A.2 Algorithm
+While prior works apply vectorized computation during the quantization step, they typically compute per-channel scaling parameters using iterative channel-wise operations.
+GranQ differs by also vectorizing the scaling computation, allowing both steps to be processed in parallel.
+Algorithm 1 outlines the complete process of quantizing and dequantizing activation tensors using GranQ.
+In Algorithm 1, both the scaling and quantization steps are implemented using vectorized tensor operations across all channels.
+Specifically, the per-channel minimum and maximum values are computed simultaneously, and the resulting scaling factors and zero-points are applied through broadcasting.
+This eliminates the need for channel-wise loops and allows efficient execution on modern hardware accelerators.
+In contrast, Algorithm 2 follows a conventional channel-wise quantization approach, where the scaling parameters are computed sequentially for each channel using an explicit for-loop.
+The quantization and dequantization steps are still vectorized per channel.
+However, the scaling computation remains scalar and iterative.
+This leads to increased computational overhead, especially when applied to deep networks with many channels.
+By removing this bottleneck, GranQ achieves both higher parallelism and lower latency, which are particularly beneficial in low-bitwidth quantization scenarios.
+Algorithm 1 GranQ: Vectorized Per-Channel Scaling for Activation Quantization
+1:Activation tensors $\{A_{l}\}_{l=1}^{L}$, where $A_{l}\in\mathbb{R}^{C\times H\times W}$, quantization bit $b$
+2:Dequantized activation tensors $\{A_{\text{deq},l}\}_{l=1}^{L}$
+3:for each layer $l$ in $L$ do
+4:    Reshape:
+$A_{l}\leftarrow\text{reshape}(A_{l})\in\mathbb{R}^{C\times(HW)}$
+$\triangleright$ Flatten spatial dimensions
+5:    Decompose by channel:
+$\vec{A}_{\min}=\min_{h,w}(A_{l}),\quad\vec{A}_{\max}=\max_{h,w}(A_{l})$
+$\triangleright$ Extract per-channel range
+6:    Compute scaling parameters:
+$\vec{s}=\frac{\vec{A}_{\max}-\vec{A}_{\min}}{2^{b}-1},\quad\vec{z}=\left\lfloor-\frac{\vec{A}_{\min}}{\vec{s}}\right\rceil$
+7:    Quantize (vectorized):
+$A_{q,l}=\left\lfloor\frac{A_{l}}{\vec{s}}+\vec{z}\right\rceil$
+8:    Dequantize (vectorized):
+$A_{\text{deq},l}=\vec{s}(A_{q,l}-\vec{z})$
+9:end for
+10:Return: $\{A_{\text{deq},l}\}_{l=1}^{L}$
+Algorithm 2 Conventional Channel-Wise Activation Quantization (Scalar-loop for scaling)
+1:Activation tensors $\{A_{l}\}_{l=1}^{L}$, where $A_{l}\in\mathbb{R}^{C\times H\times W}$, quantization bit $b$
+2:Dequantized activation tensors $\{A_{\text{deq},l}\}_{l=1}^{L}$
+3:for each layer $l$ in $L$ do
+4:    Compute per-channel scaling parameters:
+5:    for each channel $c$ in $C$ do
+$s_{c}=\frac{\max(A_{l,c})-\min(A_{l,c})}{2^{b}-1},\quad z_{c}=\left\lfloor-\frac{\min(A_{l,c})}{s_{c}}\right\rceil$
+6:    end for
+7:    Apply quantization (vectorized):
+$A_{q,l}[c,:,:]=\left\lfloor\frac{A_{l}[c,:,:]}{s_{c}}+z_{c}\right\rceil,\quad\forall c$
+8:    Dequantize (vectorized):
+$A_{\text{deq},l}[c,:,:]=s_{c}(A_{q,l}[c,:,:]-z_{c}),\quad\forall c$
+9:end for
+10:Return: $\{A_{\text{deq},l}\}_{l=1}^{L}$
+A.3 Details of Experimental Setup
+We describe the datasets, experimental setup, and baselines to ensure reproducibility and highlight both the effectiveness and implementation accessibility of our approach.
+A.3.1 Datasets
 This paper outlines the experimental settings for CIFAR-10 [13], CIFAR-100 [13], and ImageNet (ILSVRC-2012) [14] datasets.
@@
 We employ these widely used benchmarks for zero-shot quantization (ZSQ) to assess and compare the performance of GranQ against those of existing approaches.
-B.2 Experimental Setup
+A.3.2 Experimental Setup
 The data generation process followed the settings of AdaDFQ [1].
@@
 The batch size was set to 16 for CIFAR-10 and ImageNet, and 200 for CIFAR-100.
-All experiments used Nesterov-based SGD [11] as the optimizer, with a momentum of 0.9 and a weight decay of 1e-4.
+All experiments used Nesterov-based SGD [47] as the optimizer, with momentum 0.9 and weight decay 1e-4.
 In CIFAR experiments, the weight decay of 1e-4 was reduced by a factor of 0.1 every 100 epochs, whereas for ImageNet, decay was applied at 350 and 400 epochs (for ResNet-50 at 150, 300, and 400 epochs).
@@
 All experiments were conducted on an NVIDIA RTX 3090 GPU.
-B.3 Comparison Methods
+A.3.3 Comparison Methods
 GranQ is a quantization-aware training (QAT) method for zero-shot quantization (ZSQ).
 Therefore, to evaluate the proposed approach, we conducted comparative experiments using the following state-of-the-art methods:
-Algorithm 1 GranQ: Layer-Channel Quantization
-1:Activation tensors $\{x_{l}\}_{l=1}^{L}$, where $x_{l}\in\mathbb{R}^{C\times(HW)}$, quantization bit $b$
-2:Dequantized activation tensors $\{x_{\text{deq},l}\}_{l=1}^{L}$
-3:for each layer $l$ in $L$ do $\triangleright$ Process each layer separately
-4:       Compute per-channel min/max values:
-$\vec{x}_{\min}=\min_{h,w}(x_{l}),\quad\vec{x}_{\max}=\max_{h,w}(x_{l})$
-$\triangleright$ Compute along the channel axis for layer $l$
-5:       Compute scaling factors and zero-points (Vectorized):
-$\vec{s}=\frac{\vec{x}_{\max}-\vec{x}_{\min}}{2^{b}-1},\quad\vec{z}=\left\lfloor-\frac{\vec{x}_{\min}}{\vec{s}}\right\rceil$
-6:       Quantize values (Vectorized):
-$x_{q,l}=\left\lfloor\frac{x_{l}}{\vec{s}}+\vec{z}\right\rceil$
-$\triangleright$ Element-wise quantization for all channels in layer $l$
-7:       Dequantize values (Vectorized):
-$x_{\text{deq},l}=\vec{s}(x_{q,l}-\vec{z})$
-$\triangleright$ Reconstruct activation tensor for layer $l$
-8:end for
-9:Return the dequantized tensors:
-$\textbf{return}\ \{x_{\text{deq},l}\}_{l=1}^{L}$
   • GDFQ [5]: GDFQ (Generative low-bit data-free quantization) is the first approach to train quantized models using synthetic data through knowledge distillation.
@@
   • RIS [8]: RIS (Robustness-guided image synthesis) addressed the low semantic quality of generated data and proposed using perturbations to improve data diversity.
-  • PLF [5]: PLF was the first to evaluate generated data before quantization, offering a method to ensure more reliable data.
   • GenQ [10]: GenQ introduced the use of stable diffusion in ZSQ, enabling the generation of high-resolution, realistic data.
@@
   • SynQ [12]: SynQ enhanced the training of quantized models by reducing noise in generated data using low-pass filtering and aligning activation maps.
-Figure 6: Ablation study on relative error in ResNet-20 across layers.
-Layer-Q, Channel-Q, and GranQ represent layer-wise, channel-wise, and proposed quantization.
-Experiments used AdaDFQ on CIFAR-100 with 3-bit quantization.
-Appendix C Algorithm
-GranQ is a granular quantization method that processes activation values across all channels within each layer.
-Algorithm 1 describes the procedure for quantizing the activation tensor $\{x_{deq,l}\}^{L}_{l=1}$.
-The activation tensor of layer $l$ has a shape of $C\times(HW)$, and the minimum and maximum values are calculated independently for each channel.
-Based on these values, a scaling factor and zero-point are computed per channel for the quantization process.
-The resulting quantized values are dequantized back to floating-point representation prior to returning.
-Appendix D Extra Analysis
-D.1 Analysis of Activation Preservation
-To analyze the severe quantization error (activation distortion) occurring in specific layers, we conducted additional experiments, as illustrated in Figure 6.
-This figure presents the relative error for each layer of ResNet-20 and reveals that certain layers such as the 18th exhibit significantly high errors.
-These findings suggest that both layer- and channel-wise quantization fail to effectively preserve activations in specific layers.
-In contrast, GranQ achieves consistently low errors across all layers, ensuring better activation preservation.
-D.2 Analysis of Activation Magnitude
-In Figure 7, we observe that channel-wise quantization is ineffective in ZSQ settings, similar to layer-wise quantization.
-As shown in Figure 7(a), the full-precision (FP) model exhibits significant activation magnitude differences across layers, even within the same channel axis.
-This highlights the limitations of channel-wise quantization in handling quantization errors.
-Conversely, GranQ addresses this issue by accounting for each channel across all layers, thereby minimizing quantization errors.
-This is particularly important in the final layers, where maintaining the activations of the FP model is critical for accurate inference.
-Whereas (a) fails to preserve these distributions, (b) successfully maintains them.
-(a) Layer Quantization
+These methods cover various aspects of ZSQ, including data generation, training objectives.
+We compare them to demonstrate the effectiveness and generality of GranQ, which focuses on improving the quantization process itself by introducing an optimized, vectorized scaling mechanism.
+This structural enhancement allows GranQ to be seamlessly integrated into existing QAT-based ZSQ pipelines, providing consistent gains across different settings.
+Furthermore, as several methods have not yet released official code, we expect that GranQ can be readily applied as their implementations become publicly available.
+A.4 Extra Analysis
+A.4.1 Data Distribution Analysis
+(a) Original inputs (Skewness: 0.2599)
+(b) Synthetic inputs (Skewness: 0.4877)
+Figure 6: Pixel distribution comparison between (a) original and (b) synthetic inputs on ImageNet.
+The synthetic inputs, generated by AdaDFQ [17], exhibit higher skewness and a rightward shift, indicating the presence of more extreme activation values and motivating the need for granular scaling in quantization.
+Figure 6 compares the pixel distributions of original and synthetic inputs derived from ImageNet.
+The skewness of the original input is 0.2599, whereas that of the synthetic input increases to 0.4877, with the distribution shifted toward higher pixel values.
+This indicates that synthetic data, often generated without real image priors, tends to contain a greater number of extreme activation values.
+As a result, the likelihood of outliers increases, which fixed-range quantization cannot effectively handle.
+In particular, when a single global range is applied across all channels, the presence of outliers can compress the majority of values into a narrow dynamic range, leading to severe quantization distortion.
+These observations underscore the limitations of conventional scaling strategies.
+To address this, we propose a granular, channel-wise scaling approach that dynamically adapts to per-channel statistics and better preserves activation fidelity under distributional mismatch.
+A.4.2 Computational Efficiency Analysis
+Although GranQ and conventional channel-wise quantization share a higher theoretical complexity of $\mathcal{O}(L\cdot C\cdot H\cdot W)$ compared to the layer-wise baseline with $\mathcal{O}(L\cdot H\cdot W)$, all three methods apply their respective scaling strategies only during the training phase.
+Since QAT fixes the scaling factors as static parameters prior to deployment, these operations are not executed during inference, resulting in no runtime overhead.
+Consequently, the per-channel granularity used in GranQ does not incur any additional inference latency.
+As shown in Table 7, all methods achieve comparable inference latency regardless of the scaling strategy, demonstrating that GranQ maintains runtime efficiency while benefiting from fine-grained scaling during training.
+Table 7: Comparison of activation quantization latency on ResNet-18 (ImageNet, batch size = 128) under a 3-bit quantization setting.
+All latency values are measured per batch and averaged over 3 evaluations, with standard deviation reported in parentheses.
+Theoretical complexity reflects quantization-time operations and does not affect inference latency, which is measured separately.
+Method Theoretical Complexity Scaling Strategy Inference Latency (ms)
+Activation (Layer-wise) $O(L\cdot H\cdot W)$ Layer-wise uniform scaling 44.194 (+0.190)
+Activation (Channel-wise, Scalar-loop) $O(L\cdot C\cdot H\cdot W)$ Channel-wise scalar scaling (loop-based) 44.272 (+0.158)
+GranQ (Channel-wise, Vectorization) $O(L\cdot C\cdot H\cdot W)$ Channel-wise vectorized scaling 44.231 (+0.186)
+A.4.3 Analysis of Activation Preservation
+We analyzed how activation distortion propagates across layers in a ZSQ setting where synthetic data is used as input.
+When coarse quantization is applied without sufficiently granular scaling, distortion tends to accumulate in specific layers and ultimately degrades the output layer’s activation representation.
+This suggests that insufficient preservation of intermediate activations can negatively affect the quality of final inference.
+The problem becomes more severe in zero-shot scenarios due to the distributional mismatch introduced by synthetic data.
+In contrast, GranQ mitigates this issue by applying fine-grained channel-wise scaling, resulting in consistently stable activation distributions throughout the network.
+This phenomenon is clearly demonstrated in Figure 7, where GranQ maintains stable activations with noticeably reduced distortion compared to conventional quantization methods.
+(a) Layer-wise Quantization
 (b) GranQ (Ours)

### v5

@@
 However, these synthetic inputs often lead to activation distortion, especially under low-bit settings.
-As a result, existing methods struggle to mitigate this issue due to coarse activation scaling.
-To address this issue, we propose GranQ, a novel activation quantization framework that efficiently applies per-channel scaling through vectorized computation.
-In contrast to conventional channel-wise methods, which apply vectorization only to the quantization step, GranQ improves efficiency by vectorizing the scaling operation.
-This design allows GranQ to maintain fine-grained quantization granularity with minimal computational overhead, even in low-bit environments.
+To mitigate this, existing methods typically employ per-channel scaling, but they still struggle due to the severe computational overhead during the accumulation process.
+To overcome this critical bottleneck, we propose GranQ, a novel activation quantization framework that introduces an efficient pre-scaling strategy.
+Unlike conventional channel-wise methods that repeatedly perform scaling operations during accumulation, GranQ applies scaling factors in a pre-scaling step through fully vectorized computation, eliminating runtime scaling overhead.
+This design enables GranQ to maintain fine-grained quantization accuracy while significantly reducing computational burden, particularly in low-bit quantization settings.
 Extensive experiments under quantization-aware training (QAT) settings demonstrate that GranQ consistently outperforms state-of-the-art ZSQ methods across CIFAR and ImageNet.
@@
 With these findings, we anticipate that GranQ will inspire future research beyond conventional ZSQ approaches centered on data generation and model fine-tuning.
-1 Introduction
+The official code is available at https://github.com/anonymus-orange/GranQ.
+Introduction
 Neural network compression has been extensively studied for the practical deployment of large-scale deep learning (DL) models.
 In particular, reducing model size while minimizing performance degradation is crucial for utilizing DL on edge devices (e.g., mobile phones, embedded systems, and drones).
-Major approaches to model compression include quantization [1, 2, 3], pruning [4, 5, 6, 7], knowledge distillation [8, 9, 10], and neural architecture search [11, 12], as surveyed in [13, 14].
+Major approaches to model compression include quantization (Hubara et al.
+2018; Li et al.
+2021; Gholami et al.
+2022), pruning (LeCun, Denker, and Solla 1989; Frankle and Carbin 2018; Liu et al.
+2018; Cheng, Zhang, and Shi 2024), knowledge distillation (Hinton 2015; Heo et al.
+2019; Wang and Yoon 2021), and neural architecture search (Zoph 2016; Elsken, Metzen, and Hutter 2019), as surveyed in (Cheng et al.
+2018; Deng et al.
+2020).
 Among these, quantization has emerged as the most actively studied technique.
 It serves as an effective compression method by reducing unnecessary representational ranges in the model.
-However, it often requires fine-tuning or calibration to match full-precision (FP) model performance [3].
+However, it often requires fine-tuning or calibration to match full-precision (FP) model performance (Gholami et al.
+2022).
 To address this, zero-shot quantization (ZSQ), also known as data-free quantization, has been proposed to compress models without the original training data.
-Since the introduction of ZeroQ [15], studies on ZSQ have advanced in two main directions.
+Data Generation
+(PTQ, QAT)
+Calibration
+Fine-tuning
+ZeroQ (CVPR 20)
+GDFQ (ECCV 20)
+DSG (CVPR 21)
+MixMix (CVPR 21)
+Qimera (NeurIPS 21)
+IntraQ (CVPR 22)
+GENIE (CVPR 23)
+AdaDFQ (CVPR 23)
+Casual-DFQ (ICCV 23)
+TexQ (NeurIPS 23)
+RIS (AAAI 24)
+GenQ (ECCV 24)
+SQuant (ICLR 22)
+UDFC (ICCV 23)
+AIT (CVPR 22)
+PLF (CVPR 24)
+SynQ (ICLR 25)
+GranQ (Ours)
+Table 1: Categorizing ZSQ algorithms.
+PTQ denotes post-training quantization, and QAT denotes quantization-aware training.
+Data generation methods typically support both PTQ and QAT, whereas violet-marked methods utilize fine-tuning for data generation and support only QAT.
+Since the introduction of ZeroQ (Cai et al.
+2020), studies on ZSQ have advanced in two main directions.
 The first direction focuses on data generation, where synthetic data are created from the FP model.
@@
 We categorize existing ZSQ methods based on these directions, as summarized in Table 1.
-First, data generation studies focus on generating high-quality data to effectively train the Q model [16, 17, 18].
-Meanwhile, calibration (PTQ) studies aim to minimize the quantization error by calibrating the Q model without additional training [19, 20].
-Finally, fine-tuning (QAT) studies focus on transferring key information from the FP model to the Q model through knowledge distillation [21, 22, 23, 24].
-However, despite extensive studies, severe performance degradation in low-bit quantization remains unresolved.
-To address this issue, we performed an in-depth analysis of the ZSQ process, focusing on why low-bitwidth settings still suffer from performance loss even after QAT fine-tuning.
-Our findings reveal that quantization errors mainly stem from the loss of activation values instead of data quality or training methods.
-Notably, we found that layer-wise (per-tensor) quantization is no longer suitable for preserving activations in ZSQ, as it leads to coarse and inaccurate representations.
-Based on this analysis, we introduce GranQ, a novel ZSQ approach that enables channel-wise activation quantization by adapting the scale to each channel of every input activation.
-This allows us to overcome the limitations of conventional activation quantization.
-This dynamic adjustment minimizes activation loss and preserves the original activation values by reducing quantization errors, as shown in Figure 1.
-The proposed method effectively handles activation loss in low-bit quantization and achieves state-of-the-art (SOTA) performance in QAT settings on the CIFAR and ImageNet datasets.
-Furthermore, We apply vectorization to the scaling operation, which is not applied in conventional channel-wise quantization, reducing latency while enabling fine-grained activation quantization.
-Our contributions can be summarized as follows:
 (a) Layer-wise Quantization
@@
 GranQ better preserves the original activation with minimal distortion.
+First, data generation studies focus on generating high-quality data to effectively train the Q model (Xu et al.
+2020; Qian et al.
+2023; Li et al.
+2024).
+Meanwhile, calibration (PTQ) studies aim to minimize the quantization error by calibrating the Q model without additional training (Guo et al.
+2022; Bai et al.
+2023).
+Finally, fine-tuning (QAT) studies focus on transferring key information from the FP model to the Q model through knowledge distillation (Choi et al.
+2022; Fan et al.
+2024; Hong et al.
+2025; Kim, Kim, and Kang 2025).
+However, despite extensive studies, severe performance degradation in low-bit quantization remains unresolved.
+To address this issue, we performed an in-depth analysis of the ZSQ process, focusing on why low-bitwidth settings still suffer from performance loss even after QAT fine-tuning.
+Our findings reveal that quantization errors mainly stem from the loss of activation values instead of data quality or training methods.
+Notably, we found that layer-wise (per-tensor) quantization is no longer suitable for preserving activations in ZSQ, as it leads to coarse and inaccurate representations.
+Based on this analysis, we introduce GranQ, a novel ZSQ method that achieves efficient per-channel quantization via vectorized pre-scaling of input-dependent activations.
+This dynamic adjustment minimizes activation loss and preserves the original activation values by reducing quantization errors, as shown in Figure 1.
+The proposed method effectively handles activation loss in low-bit quantization and achieves state-of-the-art (SOTA) performance in QAT settings on the CIFAR and ImageNet datasets.
+Furthermore, we apply vectorization to the pre-scaling step, which is typically omitted in conventional channel-wise quantization, thereby reducing latency and enabling fine-grained activation quantization.
+Our contributions can be summarized as follows:
   • We identify critical limitations of layer-wise activation quantization in low-bit ZSQ.
@@
   • We propose GranQ, a novel method that supports granular quantization and maintains computational efficiency.
-We introduce a fine-grained activation quantization mechanism that considers every channel to reduce activation loss.
 Although per-channel activation quantization is known to improve precision, it has not been widely adopted due to its high computational cost.
-We address this by applying vectorization to the scaling operation, enabling efficient per-channel quantization with minimal overhead.
+We address this by introducing vectorized pre-scaling, which integrates per-channel scaling into the quantization step, allowing accumulation to proceed efficiently without runtime scaling overhead.
 To the best of our knowledge, our approach is the first to address the ZSQ problem.
@@
 Furthermore, on the CIFAR-10 dataset in the 5-bit quantization setting, GranQ achieves an accuracy of 94.06%, slightly exceeding the FP model performance by 0.17%.
-2 Related Work
-2.1 Quantization
+Related Work
+Quantization
 Quantization reduces the representational range of deep neural networks (DNNs) and minimizes memory usage.
-It is typically categorized into post-training quantization (PTQ) and quantization-aware training (QAT) [3, 25].
+It is typically categorized into post-training quantization (PTQ) and quantization-aware training (QAT) (Gholami et al.
+2022; Banner, Nahshan, and Soudry 2019).
 PTQ applies quantization after training without further updates, which makes it efficient and easy to implement.
 However, it is sensitive to scaling errors and often relies on calibration to improve accuracy.
-In contrast, QAT incorporates quantization during training and optimizes quantized activations using the straight-through estimator [3, 26].
-Both methods require data—PTQ for calibration and QAT for fine-tuning.
+In contrast, QAT incorporates quantization during training and optimizes quantized activations using the straight-through estimator (Gholami et al.
+2022; Yin et al.
+2019).
+Both methods require data.
+PTQ uses it for calibration, while QAT uses it for fine-tuning..
 In practice, access to training data is often limited.
-To address this, zero-shot quantization (ZSQ) has been proposed to perform quantization without using original data [16].
-Data Generation
-(PTQ, QAT)
-Calibration
-Fine-tuning
-ZeroQ (CVPR 20) [15]
-GDFQ (ECCV 20) [16]
-DSG (CVPR 21) [27]
-MixMix (CVPR 21) [28]
-Qimera (NeurIPS 21) [29]
-IntraQ (CVPR 22) [30]
-GENIE (CVPR 23) [31]
-AdaDFQ (CVPR 23) [17]
-Casual-DFQ (ICCV 23) [32]
-TexQ (NeurIPS 23) [33]
-RIS (AAAI 24) [34]
-GenQ (ECCV 24) [18]
-SQuant (ICLR 22) [19]
-UDFC (ICCV 23) [20]
-AIT (CVPR 22) [21]
-PLF (CVPR 24) [22]
-SynQ (ICLR 25) [24]
-GranQ (Ours)
-Table 1: Categorizing zero-shot quantization algorithms.
-PTQ denotes post-training quantization, and QAT denotes quantization-aware training.
-Data generation methods typically support both PTQ and QAT, whereas violet-marked methods utilize fine-tuning for data generation and support only QAT.
-2.2 Zero-shot Quantization
-As summarized in Table 1, ZSQ was initially introduced by ZeroQ [15], leading to the development of various ZSQ algorithms.
+To address this, zero-shot quantization (ZSQ) has been proposed to perform quantization without using original data (Xu et al.
+2020).
+Zero-shot Quantization
+As summarized in Table 1, ZSQ was initially introduced by ZeroQ (Cai et al.
+2020), leading to the development of various ZSQ algorithms.
 These studies have explored diverse approaches to improve quantization performance under data-free settings.
-Recent studies have focused on generating high-quality synthetic data [31, 18] and developing more effective methods to utilize them [21, 23, 24].
+Recent studies have focused on generating high-quality synthetic data (Jeon, Lee, and Kim 2023; Li et al.
+2024) and developing more effective methods to utilize them (Choi et al.
+2022; Hong et al.
+2025; Kim, Kim, and Kang 2025).
 However, current ZSQ methods heavily rely on augmented synthetic inputs.
@@
 In practice, such input-dependent channel-wise quantization is rarely adopted due to its high computational cost.
-Therefore, recent studies have applied channel-wise activation quantization either in a limited conditions within QAT settings [35], or exclusively in PTQ scenarios with access to a small amount of data [36, 37].
-This highlights the need for efficient channel-wise activation quantization for QAT in ZSQ.
-3 Preliminaries and Problem Definition
+Therefore, recent studies have applied channel-wise activation quantization either in a limited conditions within QAT settings (Cho and Yoo 2020), or exclusively in PTQ scenarios with access to a small amount of data (Yvinec et al.
+2023; Xu et al.
+2024).
+This highlights the importance of efficient channel-wise activation quantization in QAT.
+Preliminaries and Problem Definition
 In this section, we provide a new definition of the existing ZSQ problem and introduce the preliminaries. The detailed mathematical notations are listed in Table 6 in the appendix.
-3.1 Activation Quantization
-Activation quantization reduces the precision of intermediate activation values by converting them into low-bitwidth integers [38].
+Activation Quantization
+Activation quantization reduces the precision of intermediate activation values by converting them into low-bitwidth integers (Choi et al.
+2018).
 It is commonly used with weight quantization to compress models, and the linear quantization scheme is most widely adopted.
-This process typically consists of quantization, scaling parameter computation, and dequantization [1, 3, 39].
+This process typically consists of quantization, scaling parameter computation, and dequantization (Hubara et al.
+2018; Gholami et al.
+2022; Fang et al.
+2020).
 The quantization operator $\mathcal{Q}$ maps a floating-point value $x$ into an integer $x_{q}$ using a scaling factor $s$ and zero-point $z$:
@@
 where $x_{\min}$ and $x_{\max}$ denote the range of activation values, and $b$ is the bit-width.
-A large range yields a large $s$, increasing quantization error; a smaller range allows finer resolution.
-During inference, the dequantization step restores quantized values to real numbers:
-$x_{\text{deq}}=s(x_{q}-z)$ (3)
-This enables recovery of activation information needed for final predictions.
-3.2 Problem Definition
+A large range yields a large $s$, increasing quantization error, a smaller range allows finer resolution.
+As shown below, in quantized models, per-channel scaling is embedded directly into the accumulation path, requiring each channel to be individually scaled during computation:
+$y_{l}=\sum_{c=1}^{C}w_{c}\cdot s_{c}\cdot(x_{q,c}-z_{c})$ (3)
+However, applying the per-channel scaling factor $s_{c}$ inside the accumulation loop introduces computational overhead, making efficient integer-domain execution difficult.
+This scaling bottleneck hinders parallelism in core operations such as convolution and matrix multiplication, reducing the efficiency of quantized models (Nagel et al.
+2021).
+Problem Definition
 Existing activation quantization methods are designed based on layer-wise quantization to minimize computational cost.
@@
 These synthetic distributions are often biased or irregular, making precise quantization more difficult.
-(Problem 2) Severe Quantization Error in Certain Layers
-During activation quantization, certain layers induce severe quantization errors, leading to significant activation distortions.
-These distortions substantially differ from the original distribution and can directly impact the performance of the model.
-4 Observation and Methodology
-We observe the causes of activation distortion during quantization and analyze how to mitigate them. Based on this, we introduce the proposed method known as GranQ.
-Figure 2: Activation magnitudes of ResNet-20 on CIFAR-10 using 32-bit FP (left) and 3-bit layer-wise quantization (right). Severe distortions appear, especially in the final layer.
-4.1 Observation
-Layer-wise quantization effectively captures the activation distribution of each layer.
-However, it becomes inefficient when activation distributions within a layer exhibit significant variation along the channel axis.
-Particularly in ZSQ, as shown in the FP graph in Figure 1(a), the activation distributions can differ substantially across channels even within the same layer.
-This implies that limited representation bits are forced to capture a wide range of activation values, making fine-grained quantization difficult.
-It consistently appears across all layers.
-We analyzed this phenomenon occurring across all layers and provided a summary in Table 2.
-It presents the cosine similarity $\frac{X\cdot Q}{|X|_{2}|Q|_{2}}$, along with the relative error $\frac{|X-Q|_{2}}{|X|_{2}}$.
-From the results, we observe that layer-wise quantization fails to preserve activation information effectively.
-This reveals that using a single scaling range for activation quantization has a critical limitation, particularly in ZSQ.
-In contrast, GranQ, which applies scaling per channel, achieves 1.34$\times$ higher similarity and 2.94$\times$ lower relative error than layer-wise quantization, demonstrating its effectiveness in preserving activation information.
-Furthermore, Figure 2 provides a visual representation of the effect of quantization on activation values across layers.
-In layer-wise quantizaiton, although the overall activations generally reflect the FP activations to some extent, certain layers exhibit fine activation values being mapped to discrete values, leading to coarse representations.
-These coarse representations have been identified as a major factor leading to a decline in the overall expressiveness and inference performance of the model.
-Further analysis of Figure 2 is provided in Figure 5 and in Appendix (Figure 7).
+(Problem 2) Accumulation Bottleneck in Channel-wise Quantization The main challenge in channel-wise quantization lies not in quantizing activations but in the repeated per-channel scaling during accumulation.
+This disrupts vectorized execution and significantly increases computational overhead.
+Observation and Methodology
+We observe the causes of activation distortion during quantization and analyze how to mitigate them. Based on this, we introduce our method, GranQ.
+Observation
 Activaiton Quantization Layer-wise Channel-wise (GranQ)
@@
 Table 2: Average cosine similarity and relative error for 3-bit activation quantization on ResNet-20 (CIFAR-100).
-GranQ (channel-wise) outperforms layer-wise quantization by achieving less distortion in activation representation, as indicated by higher similarity and lower error.
-4.2 Methodology
-Based on our analysis, we propose GranQ, a fine-grained ZSQ method designed to reduce computational overhead (Figure 3).
-GranQ applies channel-wise scaling through activation reshaping and decomposition, and achieves fast parallel processing via vectorization of scaling parameters, which enables efficient computation of both scaling and quantization.
-4.2.1 Activation Reshaping and Decomposition
-We propose activation reshaping and decomposition, which reshapes each activation map from a three-dimensional tensor of shape $(C\times H\times W)$ into a two-dimensional matrix of shape $(C\times HW)$.
+GranQ shows lower activation distortion than layer-wise, with higher similarity and lower error.
+[…]
+[…]
+Furthermore, Figure 1(a) illustrates how layer-wise quantization leads to coarse activation representations across layers.
+Fine-grained activation values are often mapped to limited discrete levels, resulting in a loss of detail and reduced representational capacity.
+This degradation contributes to decreased inference performance.
+Methodology
+Based on our analysis, we propose GranQ, a fine-grained ZSQ method designed to reduce computational overhead, as illustrated in Figure 2.
+GranQ applies channel-wise scaling through activation decomposition, and achieves fast parallel processing via pre-scaling for parallel accumulation, which enables efficient computation of both scaling and quantization.
+Algorithm 1 GranQ: Pre-scaling for Parallel Accumulation
+1:Activation tensors $\{A_{l}\}_{l=1}^{L}$, where $A_{l}\in\mathbb{R}^{C\times H\times W}$, quantization bit $b$, weight tensors $\{w_{l,c}\}_{l=1,c=1}^{L,C}$
+2:Output vectors $\{y_{l}\}_{l=1}^{L}$
+3:for each layer $l$ in $L$ do
+4:  $A_{l}\leftarrow\text{reshape}(A_{l})\in\mathbb{R}^{C\times(HW)}$
+5:  $\vec{A}_{\min},\vec{A}_{\max}\leftarrow\min_{h,w}(A_{l}),\max_{h,w}(A_{l})$
+6:  $\vec{s}_{l}\leftarrow(\vec{A}_{\max}-\vec{A}_{\min})/(2^{b}-1)$
+7:  $\vec{z}_{l}\leftarrow\left\lfloor-\vec{A}_{\min}/\vec{s}_{l}\right\rceil$
+8:  $\vec{A}_{q,l}\leftarrow\left\lfloor A_{l}/\vec{s}_{l}+\vec{z}_{l}\right\rceil$ $\triangleright$ quantize (int mapping)
+9:  $\vec{\tilde{A}}_{l,c}\leftarrow\vec{s}_{l}\cdot(\vec{A}_{q,l}-\vec{z}_{l})$ $\triangleright$ dequantize & pre-scale
+10:  Compute output:
+$y_{l}=\sum_{c=1}^{C}w_{l,c}\cdot\vec{\tilde{A}}_{l,c}$
+11:end for
+12:return $\{y_{l}\}_{l=1}^{L}$
+Activation Decomposition
+We propose activation decomposition, which reshapes each activation map from a three-dimensional tensor of shape $(C\times H\times W)$ into a two-dimensional matrix of shape $(C\times HW)$.
 This transformation enables decomposition of the activation map along the channel axis, allowing each channel to be processed independently.
 It also facilitates vectorized computation of per-channel statistics (e.g., min and max) required for scaling factor calculation during quantization.
-The activation reshaping and decomposition is defined as follows:
+We define the activation decomposition as shown in Equation 4.
 $\begin{gathered}A_{l}=\left[A_{l}(1,:,:),\,A_{l}(2,:,:),\,\dots,\,A_{l}(C,:,:)\right]\in\mathbb{R}^{C\times H\times W}\\ \vec{A}_{\min}=\left[\min_{(h,w)}A_{l}(c,h,w)\right]_{c=1}^{C}\in\mathbb{R}^{C}\\ \vec{A}_{\max}=\left[\max_{(h,w)}A_{l}(c,h,w)\right]_{c=1}^{C}\in\mathbb{R}^{C}\end{gathered}$ (4)
 […]
-4.2.2 Vectorization of Scaling Parameters
-In the vectorization of scaling parameters stage, multiple scaling ranges are stored in a vectorized form for each channel.
-This allows dynamic adaptation to the diverse activation distributions induced by synthetic inputs.
-Thus, GranQ contrasts with traditional per-channel quantization, which uses fixed scaling parameters for each channel regardless of input variation.
-As shown in Equation 5, the scaling factors $\vec{s}$ and zero-points $\vec{z}$ are computed in vectorized form.
-The activation values are then quantized and dequantized according to Equation 6, with all operations applied in parallel across channels.
+Pre-scaling for Parallel Accumulation
 $\vec{s}=\frac{\vec{A}_{\max}-\vec{A}_{\min}}{2^{b}-1},\quad\vec{z}=\left\lfloor-\frac{\vec{A}_{\min}}{\vec{s}}\right\rceil$ (5)
-$A_{q,l}=\left\lfloor\frac{A}{\vec{s}}+\vec{z}\right\rceil,A_{\text{deq},l}=\vec{s}(A_{q,l}-\vec{z})$ (6)
-This process allows efficient and adaptive quantization of activation tensors exhibiting high variation across inputs, leading to improved computational efficiency (Table 7 in appendix).
-[…]
-5 Experiments
+$y_{l}=\sum_{c=1}^{C}w_{l,c}\cdot\vec{\tilde{A_{l,c}}},\quad\text{where }\vec{\tilde{A_{l,c}}}=\vec{s}_{c}\cdot\left(\left\lfloor\frac{\vec{A}_{l,c}}{\vec{s}_{c}}+\vec{z}_{c}\right\rceil-\vec{z}_{c}\right)$ (6)
+In the pre-scaling for parallel accumulation stage, the scaling factors $\vec{s}_{c}$ and zero-points $\vec{z}_{c}$ are computed in a vectorized form across channels to dynamically adapt to the diverse activation distributions induced by synthetic inputs.
+This contrasts with conventional per-channel quantization, which applies fixed scaling parameters per channel regardless of input variation.
+As shown in Equation 6, each activation $\vec{A}_{l,c}$ is first quantized using its corresponding vectorized scaling and zero-point, and then dequantized to reconstruct the floating-point value.
+The resulting $\vec{\tilde{A_{l,c}}}$ represents a pre-scaled activation that is already adjusted for accumulation.
+By applying vectorized scaling, quantization, and dequantization uniformly across channels, this process enables efficient integer-domain computation and parallel accumulation with weights $w_{l,c}$.
+Experiments
 In this section, we thoroughly evaluate the effectiveness of GranQ. Experiments are conducted on diverse benchmark datasets, with the performance compared with those of existing ZSQ methods.
-5.1 Experimental Setup and Details
-The experiments were conducted using widely adopted ZSQ evaluation datasets, including CIFAR-10, CIFAR-100 [40], and ImageNet (ILSVRC 2012) [41] validation datasets.
-For the CIFAR datasets, ResNet-20 [42] was used as the quantization model, whereas ResNet-18 [42], ResNet-50 [42], and MobileNetV2 [43] were employed for ImageNet.
-All experiments were conducted using the SGD optimizer [44] with a momentum of 0.9 and weight decay of 1e-4.
+Experimental Setup and Details
+The experiments were conducted using widely adopted ZSQ evaluation datasets, including CIFAR-10, CIFAR-100 (Krizhevsky, Hinton et al.
+2009), and ImageNet (ILSVRC 2012) (Deng et al.
+2009) validation datasets.
+For the CIFAR datasets, ResNet-20 (He et al.
+2016) was used as the quantization model, whereas ResNet-18 (He et al.
+2016), ResNet-50 (He et al.
+2016), and MobileNetV2 (Sandler et al.
+2018) were employed for ImageNet.
+All experiments were conducted using the SGD optimizer (Ruder 2016) with a momentum of 0.9 and weight decay of 1e-4.
 The CIFAR-10 and CIFAR-100 experiments were each conducted for 200 epochs, with batch sizes of 16 and 200, respectively.
@@
 The decay steps were set to 100, 200, and 300 epochs for CIFAR, and at 350 and 400 epochs for ImageNet, with a decay rate of 0.1.
-We compared our method with existing ZSQ methods [16, 21, 17, 33, 34, 22, 18, 23, 24].
-For data generation, we followed the AdaDFQ [17] approach based on ACGAN [45].
-Further implementation details are provided in the appendix, Section A.3.2 and A.3.3.
+We compared our method with existing ZSQ methods (Xu et al.
+2020; Choi et al.
+2022; Qian et al.
+2023; Chen et al.
+2024; Bai et al.
+2024; Fan et al.
+2024; Li et al.
+2024; Hong et al.
+2025; Kim, Kim, and Kang 2025).
+For data generation, we followed the AdaDFQ (Qian et al.
+2023) approach based on ACGAN (Odena, Olah, and Shlens 2017).
+Further implementation details are provided in the appendix, Section A and A.
 In our ablation study, layer-wise quantization was applied to all layers containing activation functions, while channel-wise quantization was performed per channel at the batch level.
@@
 GranQ
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
+Cifar-10
+ResNet-20
+3ww3aa 75.11 - 84.89 86.47 - - 86.76 88.11 91.37
+4ww4aa 90.25 90.49 92.31 92.68 92.59 - 92.64 92.76 93.52
+5ww5aa 93.38 92.98 93.81 - 93.59 - 93.83 - 94.06
+Cifar-100
+ResNet-20
+3ww3aa 47.61 - 52.74 55.87 - - 54.68 57.28 62.73
+4ww4aa 63.39 61.05 66.81 67.18 65.99 - 66.94 67.34 68.79
+5ww5aa 66.12 68.40 69.93 - 69.55 - 69.75 - 70.05
+ImageNet
+ResNet-18
+3ww3aa 20.23 - 38.10 50.28 - 68.18 49.88 52.02 64.41
+4ww4aa 60.60 65.73 66.53 67.73 67.55 70.03 65.89 67.90 70.39
+5ww5aa 68.49 70.28 70.29 - 70.59 - 69.40 - 71.31
+MobileNetV2
+3ww3aa 1.46 - 28.99 32.80 - 59.15 30.56 34.21 62.42
+4ww4aa 59.43 66.47 65.41 67.07 - 69.65 64.85 67.27 70.62
+5ww5aa 68.11 71.96 71.61 - - - 71.71 - 72.49
+ResNet-50
+3ww3aa 0.31 - 17.63 25.27 - 73.99 24.50 26.89 70.76
+4ww4aa 54.16 68.27 68.38 70.72 71.54 76.10 68.75 71.05 76.63
+5ww5aa 71.63 76.00 76.03 - 76.36 - 75.90 - 77.58
 Table 3: Accuracy Evaluation of QAT Methods for ZSQ.
@@
 Baseline +GranQ Baseline +GranQ Baseline +GranQ Baseline +GranQ
-[…]
-[…]
-[…]
+3ww3aa 47.61 59.04+11.43 45.70 60.42+14.72 52.74 62.73+9.99 54.68 62.01+7.33
+4ww4aa 63.39 66.97+3.58 65.80 68.08+2.28 66.81 68.79+1.98 66.94 68.77+1.83
+5ww5aa 66.12 68.96+2.84 69.26 70.14+0.88 69.93 70.05+0.12 69.75 70.21+0.46
 Table 4: Accuracy of existing SOTA methods with the integration of GranQ. GDFQ and AdaDFQ focus on data generation, whereas Qimera+AIT and AdaDFQ+AKT primarily enhance quantized model training.
@@
 Figure 4: Relative quantization error across layers in ResNet-20 with 3-bit quantization on CIFAR-100.
-5.2 Performance Evaluation
+Performance Evaluation
 We evaluated the performance of GranQ against SOTA ZSQ methods, with results summarized in Table 3. All comparison experiments were conducted under 3, 4, and 5-bit quantization settings.
@@
 For CIFAR-100, the results were 70.05% (5-bit), 68.79% (4-bit), and 62.73% (3-bit).
-Notably, GranQ outperformed SynQ [24], the previous SOTA, by +5.45% in the CIFAR-100 3-bit setting.
+Notably, GranQ outperformed SynQ (Kim, Kim, and Kang 2025), the previous SOTA, by +5.45% in the CIFAR-100 3-bit setting.
 This result demonstrates its ability to effectively overcome the limitations of conventional low-bitwidth quantization techniques.
@@
 These results are -3.77% and -3.23% lower than those of GenQ, respectively.
-However, GenQ uses a pre-trained diffusion model [46] and prompts during data generation, leading to slower speeds compared with those of typical GAN-based methods [16].
-In contrast, GranQ improves the quantization mechanism without relying heavily on diffusion-based data generation, as GenQ does.
-As our method is adaptable to existing QAT frameworks, it can potentially achieve even higher performance when combined with GenQ’s data generation.
-5.3 Ablation Study
-5.3.1 Effectiveness Evaluation
+However, GenQ uses a pre-trained diffusion model (Ho, Jain, and Abbeel 2020) and prompts during data generation, leading to slower speeds compared with those of typical GAN-based methods (Xu et al.
+2020).
+In contrast, GranQ enhances the quantization mechanism without heavily relying on diffusion-based data generation, as GenQ does.
+Although GenQ demonstrates strong performance, its code is not publicly available, and thus we could not include it in our experiments.
+Nevertheless, since our method is compatible with existing QAT frameworks, it could potentially benefit from GenQ’s data generation if integration becomes possible.
+Ablation Study
+Effectiveness Evaluation
 GranQ consistently demonstrates performance improvements when applied to various ZSQ methods.
 As summarized in Table 4, GranQ achieves steady performance improvements when integrated with existing SOTA ZSQ techniques.
-First, we analyzed the impact of GranQ on data synthesis-based quantization methods, specifically GDFQ [16] and AdaDFQ [17].
-The integration of GranQ into GDFQ [16] led to a notable improvement, with accuracy rising from 47.61% to 59.04% (+11.43%) in the 3-bit setting and from 63.39% to 66.97% (+3.58%) in the 4-bit setting.
-Similarly, AdaDFQ [17] exhibited improvements of +9.99% in the 3-bit setting (from 52.74% to 62.73%) and +1.98% in the 4-bit setting (from 66.81% to 68.79%).
+First, we analyzed the impact of GranQ on data synthesis-based quantization methods, specifically GDFQ (Xu et al.
+2020) and AdaDFQ (Qian et al.
+2023).
+The integration of GranQ into GDFQ (Xu et al.
+2020) led to a notable improvement, with accuracy rising from 47.61% to 59.04% (+11.43%) in the 3-bit setting and from 63.39% to 66.97% (+3.58%) in the 4-bit setting.
+Similarly, AdaDFQ (Qian et al.
+2023) exhibited improvements of +9.99% in the 3-bit setting (from 52.74% to 62.73%) and +1.98% in the 4-bit setting (from 66.81% to 68.79%).
 These results suggest that GranQ effectively reduces quantization errors when combined with data synthesis-based quantization methods, leading to enhanced model performance.
-Furthermore, GranQ also exhibits significant improvements in methods focused on training quantized models, such as Qimera+AIT [21] and AdaDFQ+AKT [23].
-For Qimera+AIT [21], the 3-bit accuracy increased from 45.70% to 60.42% (+14.72%).
-Similarly, for AdaDFQ+AKT [23], the accuracy improved from 54.68% to 62.01% (+7.33%).
+Furthermore, GranQ also exhibits significant improvements in methods focused on training quantized models, such as Qimera+AIT (Choi et al.
+2022) and AdaDFQ+AKT (Hong et al.
+2025).
+For Qimera+AIT (Choi et al.
+2022), the 3-bit accuracy increased from 45.70% to 60.42% (+14.72%).
+Similarly, for AdaDFQ+AKT (Hong et al.
+2025), the accuracy improved from 54.68% to 62.01% (+7.33%).
 These findings demonstrate that GranQ is not only effective in data synthesis-based methods but also enhances performance during the model training process.
-5.3.2 Efficiency Evaluation
+Efficiency Evaluation
 Method
 Quantization
-Application
-Scaling Factor
-Computation
+Pre-Scaling
 Accuracy (%)
@@
 Latency (ms)
-Layer-wise 49.98 ($\pm$ 0.24) 0.227
-Channel-wise (Conventional scalar-loop) 62.68 ($\pm$ 0.11) 103.671
-Channel-wise (Vectorized, GranQ) 62.73 ($\pm$ 0.17) 0.236
-Table 5: Ablation study of vectorization in scaling computation and quantization.
-GranQ applies vectorization to both stages, achieving high accuracy with low latency.
-Latency is measured with batch size 16.
-Experiments use AdaDFQ on ResNet-20 (CIFAR-100, 3-bit).
+Training
+Time (sec/epoch)
+Layer-wise 49.98 0.227 10.92
+Channel-wise (Baseline) 62.68 103.671 17370.54
+Channel-wise (GranQ) 62.73 0.236 12.43
+Table 5: Ablation study on the impact of pre-scaling and vectorization in quantization.
+✓indicates whether vectorization is applied.
+Latency is measured with a batch size of 16 using AdaDFQ on ResNet-20 (CIFAR-100, 3-bit).
 While fine-grained activation quantization often incurs high computational cost due to per-channel scaling, GranQ maintains efficiency by vectorizing the scaling computation.
@@
 To assess this in practice, we measured quantization latency across various batch sizes (16, 32, 64, 128, and 200).
-Figure 5 compares three methods: layer-wise quantization, channel-wise quantization with scalar-loop scaling (Algorithm 2 in appendix), and our method (GranQ), which efficiently parallelizes scaling across channels (Algorithm 1 in appendix).
+Figure 4 compares three methods: layer-wise quantization, channel-wise quantization with scalar-loop scaling (Algorithm 2 in appendix), and our method (GranQ), which efficiently parallelizes scaling across channels.
 The two channel-wise methods perform per-channel quantization.
-The conventional method computes scaling factors sequentially, whereas GranQ vectorizes the scaling computation across channels for improved efficiency.
-A key observation from Figure 5 is that GranQ incurs only a slight increase in latency compared to conventional layer-wise quantization.
-With only a minor increase in latency, GranQ achieves higher accuracy than the layer-wise baseline, as shown in Table 5.
-Compared to scalar-loop-based channel quantization, which implements the same computation sequentially, GranQ achieves nearly identical accuracy.
-Furthermore, Table 2 shows that GranQ substantially reduces quantization error and more effectively preserves activation information.
+The conventional method computes scaling factors sequentially, whereas GranQ pre-applies them in a vectorized manner across channels, enabling pre-scaling for efficient parallel accumulation.
+A key observation from Figure 4 is that GranQ achieves substantial accuracy improvement with only a minimal latency overhead compared to conventional layer-wise quantization.
+Furthermore, Table 2 and Figure 4 shows that GranQ substantially reduces quantization error and more effectively preserves activation information.
 These results demonstrate that GranQ enables fine-grained quantization with significantly improved efficiency.
-Additional analysis and experimental results on inference efficiency are provided in Table 7 of the appendix.
-6 Discussion
-6.1 Why GranQ is Effective?
-The most common approaches to activation quantization (i.e., layer-wise quantization) are known for their low computational cost and fast processing speed.
-However, they often fail to accurately capture activation distributions, resulting in substantial quantization errors (Figure 2).
+Additionally, GranQ demonstrates that per-channel quantization can be efficiently executed by vectorizing scaling factor computation.
+As shown in Table 5, it achieves a quantization latency of 0.236 ms, comparable to the layer-wise baseline (0.227 ms), and significantly outperforms the conventional channel-wise method (103.671 ms) suffering from scalar-loop overhead.
+Discussion
+Why GranQ is Effective?
 In layer-wise quantization, representing the entire activation range with a single scaling factor makes it difficult to reflect fine-grained distribution changes.
 These limitations become more severe in low-bit settings, especially when synthetic data exhibit distributional shifts.
-To overcome these issues, GranQ computes and stores per-channel scaling factors for each input in a unified vector form, enabling parallel processing across all channels.
-Unlike conventional methods, it performs quantization based on each input’s channel-wise distribution, resulting in more precise and efficient computation.
-6.2 Limitation and Future-work
-In ZSQ settings, synthetic data often exhibit broader and more distorted activation distributions compared to original data (Figure 6 in appendix).
-GranQ addresses this challenge by applying dynamic quantization for each channel based on its activation distribution, making it particularly effective in QAT scenarios.
-However, GranQ is less suitable for PTQ, which relies on static activation quantization.
-As future work, we plan to develop quantization methods that preserve the fine-grained scaling of GranQ while enhancing adaptability to PTQ settings.
-7 Conclusion
+To overcome these issues, GranQ pre-computes per-channel scaling in a vectorized form, enabling efficient parallel accumulation.
+By integrating scaling into the quantization phase, we removed runtime scaling during accumulation, allowing precise and efficient parallel execution.
+Limitation and Future-work
+In ZSQ, synthetic data often exhibit broader and noisier activation distributions (see Figure 6 in appendix).
+While GranQ effectively handles this in QAT through dynamic per-channel scaling, its static form is less suited for PTQ.
+Future work will explore adapting GranQ to better support PTQ.
+Conclusion
 Layer-wise and channel-wise quantization are both commonly used in weight quantization, as weights remain fixed after training.
-In contrast, channel-wise scaling in activation quantization is more challenging, as activation values vary with each input, making it difficult to apply fine-grained scaling.
-These limitations become more pronounced in QAT settings, where large amounts of data are used.
-Under these conditions, quantization is strongly affected by activation distributions.
-This issue is especially critical in ZSQ, where training data are synthesized using internal statistics from the FP model.
-Synthetic data generated in this way often produce distorted activation distributions compared to original data, resulting in increased quantization error during scaling.
-To address these issues, we propose GranQ, a dynamic channel-wise quantization method that adapts scaling factors to each input’s activation distribution.
-This enables more precise quantization even under distorted distributions, as often observed in ZSQ.
-In addition, GranQ improves computational efficiency by applying per-input, per-channel scaling through vectorized operations, which is rarely adopted in activation quantization due to its implementation complexity.
-Through extensive experiments, we demonstrate that GranQ achieves superior performance while preserving computational efficiency.
-Compared with various SOTA methods, GranQ demonstrates competitive performance and achieves notable improvements.
+In contrast, it is harder to apply channel-wise activation quantization because input-dependent activations make fine-grained scaling difficult.
+These issues are especially critical in ZSQ, with data synthesized from FP statistics.
+In such scenarios, synthetic activations often deviate from real data, amplifying quantization error and complicating accurate scaling.
+In addition, channel-wise quantization poses a computational burden due to the need for per-channel scaling during accumulation.
+Since scaling factors differ across channels, accumulation requires separate multiplications, limiting parallelism and increasing latency.
+This overhead discourages the use of channel-wise methods in practice, despite their higher representational fidelity.
+To address these issues, we propose GranQ, which vectorizes scaling to enable fine-grained quantization and efficient accumulation across channels.
+Through extensive experiments, we demonstrate that GranQ achieves notable accuracy and efficiency.
 Our study highlights the importance of activation preservation in ZSQ and proposes GranQ as an insightful direction to address this challenge.
 Appendix A Appendix
-A.1 Notation
-We provide an overview of the main symbols and definitions employed throughout the paper.
-It specifically includes essential elements such as the original activation ($A$), quantized activation ($A_{q}$), scaling factor ($s$), and zero-point ($z$), along with their corresponding mathematical representations.
+Notation
+We provide an overview of the main symbols and definitions used throughout the paper, along with their precise mathematical formulations.
 Symbol Definition
-$A$ Original activation value
-$A_{q}$ Quantized activation value
-$A_{\text{deq}}$ Dequantized activation value
-$A_{\min}$, $A_{\max}$ Minimum and maximum activation values
-$Q$ Quantization operator
-$s$ Scaling factor
-$z$ Zero-point for quantization
+AA Original activation tensor
+$\vec{A}_{q}$ Quantized activation vector (per channel)
+$\vec{\tilde{A}}$ Pre-scaled activation
+$\vec{A}_{\min},\vec{A}_{\max}$ Per-channel min and max activation values
+$\vec{s}$ Per-channel scaling factor vector
+$\vec{z}$ Per-channel zero-point vector
 $b$ Number of quantization bits
-$\min\limits_{h,w}(A)$, $\max\limits_{h,w}(A)$ Channel-wise min/max activations
-$\mathbb{R}^{C\times H\times W}$ Tensor of shape $C\times H\times W$ in $\mathbb{R}$
-[…]
-$\|\cdot\|_{2}$ L2 norm (Euclidean norm)
-$\in$ Set membership ("element of")
+$C,H,W$ Channel, Height, Width dimensions of activation
+$\mathbb{R}^{C\times H\times W}$ Activation tensor space
+$\lfloor\cdot\rceil$ Rounding function
+$\cdot$ Element-wise multiplication
+$\sum_{c=1}^{C}$ Channel-wise summation
+$w_{c}$ Weight for channel $c$
+$y_{l}$ Output after accumulation at layer $l$
 Mathematical Expressions
-Forward Quantization $x_{q}=\left\lfloor\frac{x}{s}+z\right\rceil$
-Dequantization $x_{\text{deq}}=s(x_{q}-z)$
-Scaling Factor $s=\frac{x_{\max}-x_{\min}}{2^{b}-1}$
-Zero-point $z=\left\lfloor\frac{-x_{\min}}{s}\right\rceil$
+Scaling Factor $\vec{s}=\frac{\vec{A}_{\max}-\vec{A}_{\min}}{2^{b}-1}$
+Zero-point $\vec{z}=\left\lfloor\frac{-\vec{A}_{\min}}{\vec{s}}\right\rceil$
+Forward Quantization $\vec{A}_{q}=\left\lfloor\frac{A}{\vec{s}}+\vec{z}\right\rceil$
+Pre-scaled Dequantization $\vec{\tilde{A}}=\vec{s}\cdot(\vec{A}_{q}-\vec{z})$
+Output Accumulation $y=\sum_{c=1}^{C}w_{c}\cdot\vec{\tilde{A}}_{c}$
 Cosine Similarity $\frac{X\cdot Q}{\|X\|_{2}\|Q\|_{2}}$
 Relative Error $\frac{\|X-Q\|_{2}}{\|X\|_{2}}$
-Table 6: Notation for activation quantization equations, algorithm, and evaluation metrics.
-A.2 Algorithm
-While prior works apply vectorized computation during the quantization step, they typically compute per-channel scaling parameters using iterative channel-wise operations.
-GranQ differs by also vectorizing the scaling computation, allowing both steps to be processed in parallel.
-Algorithm 1 outlines the complete process of quantizing and dequantizing activation tensors using GranQ.
-In Algorithm 1, both the scaling and quantization steps are implemented using vectorized tensor operations across all channels.
-Specifically, the per-channel minimum and maximum values are computed simultaneously, and the resulting scaling factors and zero-points are applied through broadcasting.
-This eliminates the need for channel-wise loops and allows efficient execution on modern hardware accelerators.
-In contrast, Algorithm 2 follows a conventional channel-wise quantization approach, where the scaling parameters are computed sequentially for each channel using an explicit for-loop.
-The quantization and dequantization steps are still vectorized per channel.
-However, the scaling computation remains scalar and iterative.
-This leads to increased computational overhead, especially when applied to deep networks with many channels.
-By removing this bottleneck, GranQ achieves both higher parallelism and lower latency, which are particularly beneficial in low-bitwidth quantization scenarios.
-Algorithm 1 GranQ: Vectorized Per-Channel Scaling for Activation Quantization
-1:Activation tensors $\{A_{l}\}_{l=1}^{L}$, where $A_{l}\in\mathbb{R}^{C\times H\times W}$, quantization bit $b$
-2:Dequantized activation tensors $\{A_{\text{deq},l}\}_{l=1}^{L}$
-3:for each layer $l$ in $L$ do
-4:    Reshape:
-$A_{l}\leftarrow\text{reshape}(A_{l})\in\mathbb{R}^{C\times(HW)}$
-$\triangleright$ Flatten spatial dimensions
-5:    Decompose by channel:
-$\vec{A}_{\min}=\min_{h,w}(A_{l}),\quad\vec{A}_{\max}=\max_{h,w}(A_{l})$
-$\triangleright$ Extract per-channel range
-6:    Compute scaling parameters:
-$\vec{s}=\frac{\vec{A}_{\max}-\vec{A}_{\min}}{2^{b}-1},\quad\vec{z}=\left\lfloor-\frac{\vec{A}_{\min}}{\vec{s}}\right\rceil$
-7:    Quantize (vectorized):
-$A_{q,l}=\left\lfloor\frac{A_{l}}{\vec{s}}+\vec{z}\right\rceil$
-8:    Dequantize (vectorized):
-$A_{\text{deq},l}=\vec{s}(A_{q,l}-\vec{z})$
-9:end for
-10:Return: $\{A_{\text{deq},l}\}_{l=1}^{L}$
-Algorithm 2 Conventional Channel-Wise Activation Quantization (Scalar-loop for scaling)
-1:Activation tensors $\{A_{l}\}_{l=1}^{L}$, where $A_{l}\in\mathbb{R}^{C\times H\times W}$, quantization bit $b$
-2:Dequantized activation tensors $\{A_{\text{deq},l}\}_{l=1}^{L}$
-3:for each layer $l$ in $L$ do
-4:    Compute per-channel scaling parameters:
-5:    for each channel $c$ in $C$ do
-$s_{c}=\frac{\max(A_{l,c})-\min(A_{l,c})}{2^{b}-1},\quad z_{c}=\left\lfloor-\frac{\min(A_{l,c})}{s_{c}}\right\rceil$
-6:    end for
-7:    Apply quantization (vectorized):
-$A_{q,l}[c,:,:]=\left\lfloor\frac{A_{l}[c,:,:]}{s_{c}}+z_{c}\right\rceil,\quad\forall c$
-8:    Dequantize (vectorized):
-$A_{\text{deq},l}[c,:,:]=s_{c}(A_{q,l}[c,:,:]-z_{c}),\quad\forall c$
-9:end for
-10:Return: $\{A_{\text{deq},l}\}_{l=1}^{L}$
-A.3 Details of Experimental Setup
+Table 6: Notation and expressions for vectorized channel-wise activation quantization with pre-scaling.
+Details of Experimental Setup
 We describe the datasets, experimental setup, and baselines to ensure reproducibility and highlight both the effectiveness and implementation accessibility of our approach.
-A.3.1 Datasets
-This paper outlines the experimental settings for CIFAR-10 [13], CIFAR-100 [13], and ImageNet (ILSVRC-2012) [14] datasets.
+Datasets
+This paper outlines the experimental settings for CIFAR-10 (Krizhevsky, Hinton et al.
+2009), CIFAR-100 (Krizhevsky, Hinton et al.
+2009), and ImageNet (ILSVRC-2012) (Deng et al.
+2009) datasets.
 CIFAR-10 and CIFAR-100 contain 10 and 100 classes, respectively, with 32 × 32 resolution color images, and both have 50 K training and 10 K test images.
@@
 We employ these widely used benchmarks for zero-shot quantization (ZSQ) to assess and compare the performance of GranQ against those of existing approaches.
-A.3.2 Experimental Setup
-The data generation process followed the settings of AdaDFQ [1].
-Specifically, ACGAN [2] was used for data generation with the Adam optimizer, set with a momentum of 0.9 and a learning rate of 1e-3.
-All hyper-parameters such as $\alpha_{ds}$, $\alpha_{as}$, $\lambda_{l}$, and $\lambda_{u}$ followed the default settings from AdaDFQ [1].
+Experimental Setup
+The data generation process followed the settings of AdaDFQ (Qian et al.
+2023).
+Specifically, ACGAN (Odena, Olah, and Shlens 2017) was used for data generation with the Adam optimizer, set with a momentum of 0.9 and a learning rate of 1e-3.
+All hyper-parameters such as $\alpha_{ds}$, $\alpha_{as}$, $\lambda_{l}$, and $\lambda_{u}$ followed the default settings from AdaDFQ (Qian et al.
+2023).
 For the fine-tuning process, a warmup phase was applied to ensure training stability, with 4 epochs out of 200 for CIFAR and 50 epochs out of 400 for ImageNet.
 The batch size was set to 16 for CIFAR-10 and ImageNet, and 200 for CIFAR-100.
-All experiments used Nesterov-based SGD [47] as the optimizer, with momentum 0.9 and weight decay 1e-4.
+All experiments used Nesterov-based SGD (Nesterov 1983) as the optimizer, with momentum 0.9 and weight decay 1e-4.
 In CIFAR experiments, the weight decay of 1e-4 was reduced by a factor of 0.1 every 100 epochs, whereas for ImageNet, decay was applied at 350 and 400 epochs (for ResNet-50 at 150, 300, and 400 epochs).
-GranQ was implemented using PyTorch [3], and visualizations were generated with matplotlib [4].
+GranQ was implemented using PyTorch (Paszke et al.
+2019), and visualizations were generated with matplotlib (Hunter 2007).
 All experiments were conducted on an NVIDIA RTX 3090 GPU.
-A.3.3 Comparison Methods
+Comparison Methods
 GranQ is a quantization-aware training (QAT) method for zero-shot quantization (ZSQ).
 Therefore, to evaluate the proposed approach, we conducted comparative experiments using the following state-of-the-art methods:
-  • GDFQ [5]: GDFQ (Generative low-bit data-free quantization) is the first approach to train quantized models using synthetic data through knowledge distillation.
+  • GDFQ (Xu et al.
+2020): GDFQ (Generative low-bit data-free quantization) is the first approach to train quantized models using synthetic data through knowledge distillation.
 It demonstrated that effective quantization is possible even without access to original data.
-  • AIT [6]: AIT was the first to show the negative impact of cross-entropy loss when training quantized models with distillation.
+  • AIT (Choi et al.
+2022): AIT was the first to show the negative impact of cross-entropy loss when training quantized models with distillation.
 It also provided a systematic analysis of training strategies in the ZSQ domain.
-  • AdaDFQ [1]: AdaDFQ (Adaptive data-free quantization) highlighted the generalization issues of generated data under various quantization settings and proposed a more adaptable data generation method.
-  • TexQ [7]: TexQ introduced a data generation method that preserves texture distributions, thereby enhancing the quality of generated data in ZSQ environments.
-  • RIS [8]: RIS (Robustness-guided image synthesis) addressed the low semantic quality of generated data and proposed using perturbations to improve data diversity.
-  • GenQ [10]: GenQ introduced the use of stable diffusion in ZSQ, enabling the generation of high-resolution, realistic data.
-  • AKT [11]: AKT (Advanced knowledge transfer) proposed refining feature information during distillation to improve the training of quantized models.
-  • SynQ [12]: SynQ enhanced the training of quantized models by reducing noise in generated data using low-pass filtering and aligning activation maps.
-These methods cover various aspects of ZSQ, including data generation, training objectives.
-We compare them to demonstrate the effectiveness and generality of GranQ, which focuses on improving the quantization process itself by introducing an optimized, vectorized scaling mechanism.
-This structural enhancement allows GranQ to be seamlessly integrated into existing QAT-based ZSQ pipelines, providing consistent gains across different settings.
-Furthermore, as several methods have not yet released official code, we expect that GranQ can be readily applied as their implementations become publicly available.
-A.4 Extra Analysis
-A.4.1 Data Distribution Analysis
+  • AdaDFQ (Qian et al.
+2023): AdaDFQ (Adaptive data-free quantization) highlighted the generalization issues of generated data under various quantization settings and proposed a more adaptable data generation method.
+  • TexQ (Chen et al. 2023): TexQ introduced a data generation method that preserves texture distributions, thereby enhancing the quality of generated data in ZSQ environments.
+  • RIS (Bai et al. 2024): RIS (Robustness-guided image synthesis) addressed the low semantic quality of generated data and proposed using perturbations to improve data diversity.
+  • GenQ (Li et al. 2024): GenQ introduced the use of stable diffusion in ZSQ, enabling the generation of high-resolution, realistic data.
+  • AKT (Hong et al. 2024): AKT (Advanced knowledge transfer) proposed refining feature information during distillation to improve the training of quantized models.
+  • SynQ (Kim, Kim, and Kang 2025): SynQ enhanced the training of quantized models by reducing noise in generated data using low-pass filtering and aligning activation maps.
+Algorithm 2 Conventional Channel-Wise Quantization (Scalar-loop for scaling)
+1:Activation tensors $\{A_{l}\}_{l=1}^{L}$, where $A_{l}\in\mathbb{R}^{C\times H\times W}$, quantization bit $b$, weights $\{W_{l}\}_{l=1}^{L}$
+2:Output vectors $\{y_{l}\}_{l=1}^{L}$
+3:for each layer $l$ in $L$ do
+4:  $A_{l}\leftarrow\text{reshape}(A_{l})\in\mathbb{R}^{C\times(HW)}$
+5:  for each channel $c$ in $C$ do
+6:   $s_{c}=\frac{\max(A_{l,c})-\min(A_{l,c})}{2^{b}-1}$
+7:   $z_{c}=\left\lfloor-\frac{\min(A_{l,c})}{s_{c}}\right\rceil$
+8:   $A_{q,l}[c,:]=\left\lfloor\frac{A_{l}[c,:]}{s_{c}}+z_{c}\right\rceil$
+9:  end for
+10:  Compute output:
+$y_{l}=\sum_{c=1}^{C}w_{l,c}\cdot\left(s_{c}\cdot(A_{q,l}[c,:]-z_{c})\right)$
+11:end for
+12:return $\{y_{l}\}_{l=1}^{L}$
+Algorithm
+While prior works apply vectorized computation only to the quantization step, they compute per-channel scaling parameters through iterative operations.
+GranQ also vectorizes this scaling step, enabling fully parallel quantization.
+Algorithm 1 summarizes the process with pre-scaling.
+In Algorithm 1, the entire quantization procedure is implemented using vectorized tensor operations across all channels, enabling efficient accumulation without runtime scaling.
+Specifically, the per-channel minimum and maximum values are computed in parallel, and the corresponding scaling factors and zero-points are applied via broadcasting.
+This removes the need for channel-wise loops and enables efficient accumulation without runtime scaling, making the procedure highly suitable for modern hardware accelerators.
+In contrast, Algorithm 2 adopts a conventional channel-wise quantization scheme, where scaling parameters are computed iteratively for each channel using a for-loop.
+Although the quantization and dequantization steps are applied in a vectorized manner per channel, the overall process remains sequential due to scalar-based scaling.
+This incurs significant computational overhead, particularly in networks with a large number of channels.
+By removing this bottleneck, GranQ achieves both higher parallelism and lower latency, which are particularly beneficial in low-bitwidth quantization scenarios.
+Extra Analysis
+Data Distribution Analysis
 (a) Original inputs (Skewness: 0.2599)
@@
 Figure 5: Pixel distribution comparison between (a) original and (b) synthetic inputs on ImageNet.
-The synthetic inputs, generated by AdaDFQ [17], exhibit higher skewness and a rightward shift, indicating the presence of more extreme activation values and motivating the need for granular scaling in quantization.
+Synthetic inputs generated by AdaDFQ (Qian et al.
+2023) exhibit a rightward shift with higher skewness, indicating more extreme activations and highlighting the need for granular scaling.
 Figure 5 compares the pixel distributions of original and synthetic inputs derived from ImageNet.
@@
 These observations underscore the limitations of conventional scaling strategies.
-To address this, we propose a granular, channel-wise scaling approach that dynamically adapts to per-channel statistics and better preserves activation fidelity under distributional mismatch.
-A.4.2 Computational Efficiency Analysis
-Although GranQ and conventional channel-wise quantization share a higher theoretical complexity of $\mathcal{O}(L\cdot C\cdot H\cdot W)$ compared to the layer-wise baseline with $\mathcal{O}(L\cdot H\cdot W)$, all three methods apply their respective scaling strategies only during the training phase.
-Since QAT fixes the scaling factors as static parameters prior to deployment, these operations are not executed during inference, resulting in no runtime overhead.
-Consequently, the per-channel granularity used in GranQ does not incur any additional inference latency.
-As shown in Table 7, all methods achieve comparable inference latency regardless of the scaling strategy, demonstrating that GranQ maintains runtime efficiency while benefiting from fine-grained scaling during training.
-Table 7: Comparison of activation quantization latency on ResNet-18 (ImageNet, batch size = 128) under a 3-bit quantization setting.
-All latency values are measured per batch and averaged over 3 evaluations, with standard deviation reported in parentheses.
-Theoretical complexity reflects quantization-time operations and does not affect inference latency, which is measured separately.
-Method Theoretical Complexity Scaling Strategy Inference Latency (ms)
-Activation (Layer-wise) $O(L\cdot H\cdot W)$ Layer-wise uniform scaling 44.194 (+0.190)
-Activation (Channel-wise, Scalar-loop) $O(L\cdot C\cdot H\cdot W)$ Channel-wise scalar scaling (loop-based) 44.272 (+0.158)
-GranQ (Channel-wise, Vectorization) $O(L\cdot C\cdot H\cdot W)$ Channel-wise vectorized scaling 44.231 (+0.186)
-A.4.3 Analysis of Activation Preservation
+Analysis of Activation Preservation
 We analyzed how activation distortion propagates across layers in a ZSQ setting where synthetic data is used as input.